### v3

# Tight frames over the quaternions and equiangular lines

Abstract
We show that much of the theory of finite tight frames can be generalised to vector spaces over the quaternions. This includes the variational characterisation, group frames, and the characterisations of projective and unitary equivalence. We are particularly interested in sets of equiangular lines (equi-isoclinic subspaces) and the groups associated with them, and how to move them between the spaces $\mathbb{R}^{d}$, $\mathbb{C}^{d}$ and $\mathbb{H}^{d}$. We discuss what the analogue of Zauner’s conjecture for equiangular lines in $\mathbb{H}^{d}$ might be.
Key Words: finite tight frames, quaternionic equiangular lines, equi-isoclinic subspaces, equichordal subspaces, projective unitary equivalence over the quaternions, group frames, quaternionic reflection groups, representations over the quaternions, projective spherical $t$-designs, Frobenius-Schur indicator, double cover of $A_{6}$.
AMS (MOS) Subject Classifications: primary 05B30, 15B33, 20G20, 42C15, 51M20,  secondary 15B57, 20C25, 51M15, 65D30, 94A12.
1 Introduction
Tight frames are a notion of redundant orthonormal bases which is of both theoretical and practical interest [Wal18]. Their recent development has been driven by connections with algebraic combinatorics and applications to quantum physics, signal analysis and engineering. In all of these settings, tight frames for which the vectors/lines are “well spread out” are desired, with equiangular tight frames being of the most interest.
We consider tight frames over the quaternions, motivated by equiangular tight frames in $\mathbb{R}^{d}$ and $\mathbb{C}^{d}$. Given enough care, much of the theory generalises to quaternionic Hilbert space $\mathbb{H}^{d}$, including the variational characterisation, group frames and $G$-matrices, and the characterisation of projective and unitary equivalence. We consider in detail how to move between tight frames (and associated linear operators) in $\mathbb{R}^{d}$, $\mathbb{C}^{d}$ and $\mathbb{H}^{d}$.
The maximum possible number of equiangular lines in $\mathbb{R}^{d}$ is ${1\over 2}d(d+1)$, and in $\mathbb{C}^{d}$ it is $d^{2}$. The bound for real equiangular lines is rarely met, but for complex lines the bound is conjectured to hold in all cases: Zauner’s conjecture on the existence of Weyl-Heisenberg SICs [Zau10], [ACFW18]. For $\mathbb{H}^{d}$ the bound is $2d^{2}-d$, for a maximum of six equiangular lines in $\mathbb{H}^{2}$, and fifteen in $\mathbb{H}^{3}$. We give an elementary construction of five equiangular lines in $\mathbb{H}^{2}$, and investigate the maximal configuration of six equiangular lines in $\mathbb{H}^{2}$ recently obtained independently by [KF08] and [ET20]. Recently, the existence of fifteen equiangular lines in $\mathbb{H}^{3}$, viewed as a simplex in the projective space $\mathbb{H}\mathbb{P}^{2}$, has been proved by [CKM16] using a Newton-Kantorovich theorem. Based on these two data points, and my instincts (there is a lot of space in $\mathbb{H}^{d}$ and the beauty of the quaternions), I had initially thought the quaternionic version of Zauner’s conjecture:
[…]
should hold. However, calculations of [CKM16], suggests that this fails for $d=4$, and the analogous situation for the octonians is much worse. Thus, it seems that equiangular lines in $\mathbb{C}^{d}$ may be a high point for satisfying the estimates on the maximal number of equiangular lines, with real and quaternionic equiangular lines rarely meeting the bound (“filling up all the space”) due to algebraic limitations of the field involved, i.e., $\mathbb{R}$ not being algebraically closed and $\mathbb{H}$ not being commutative. Still, there is much interest in the maximal sets of equiangular lines in $\mathbb{H}^{d}$, and for those in $\mathbb{R}^{d}$ (which have been studied for over half a century). Therefore, I present the following conjecture, which can play the role of Zauner’s conjecture for the theory of quaternionic equiangular lines:
Conjecture 1
There exists more than $d^{2}$ tight equiangular lines in $\mathbb{H}^{d}$, for each $d\geq 2$.
We observe that:
  • This says “there is something going on”, i.e., there are interesting equiangular lines in $\mathbb{H}^{d}$ (ones which cannot be viewed as lines in $\mathbb{C}^{d}$) for every dimension $d$.
  • This conjecture is known to hold only for $d=2,3$, and is otherwise open.
  • For some $d$, there do exist sets of $\leq d^{2}$ tight quaternionic equiangular lines, e.g., five equiangular lines in $\mathbb{H}^{3}$ (Example 2.4) and six in $\mathbb{H}^{4}$ (Example 2.5).
  • It is conjectured in [CKM16] (Conjecture 4.2) that asymptotically there exists $N\geq(4-\sqrt{2})d$ tight equiangular lines in $\mathbb{H}^{d}$, as $d\to\infty$.
We now give the basic theory of inner product spaces over the quaternions, to a point where we are able to define and discuss tight frames over $\mathbb{H}$.
1.1 Inner products over the quaternions
The reader is assumed to be familiar with the quaternions $\mathbb{H}$ which are an extension of the complex numbers $x+iy$ to a noncommutative associative algebra over the real numbers (skew field) consisting of elements:
$q=q_{1}+q_{2}i+q_{3}j+q_{4}k=(q_{1}+q_{2}i)+(q_{3}+q_{4}i)j\in\mathbb{H},\qquad q_{j}\in\mathbb{R},$
with the (noncommutative) multiplication given by Hamilton’s famous formula that $i^{2}=j^{2}=k^{2}=ijk=-1$.
Since the multiplication is not commutative, we must distinguish between left and right vector spaces (modules) over $\mathbb{H}$. Since we wish to appropriate much of matrix theory, we take our vector spaces to be right $\mathbb{H}$-vector spaces. Thus $\mathbb{H}$-linear maps $L$ have the form
$L(v_{1}\alpha_{1}+\cdots+v_{n}\alpha_{n})=L(v_{1})\alpha_{1}+\cdots+L(v_{n})\alpha_{n},$
and can be represented by matrices, with the usual rules for multiplication, i.e.,
$(AB)_{jk}=\sum_{\ell}a_{j\ell}b_{\ell k},$
where order of multiplication in $a_{j\ell}b_{\ell k}$ cannot be reversed (see [Zha97]). For those who may have noticed, I apologise for using $j$ and $k$ above as indices for matrix entries, and elsewhere as quaternion units (as is often done with the complex unit $i$).
The conjugate and norm of a quaternion $q=q_{1}+q_{2}i+q_{3}j+q_{4}k\in\mathbb{H}$
$\overline{q}:=q_{1}-q_{2}i-q_{3}j-q_{4}k,\qquad|q|:=\sqrt{q\overline{q}}=\sqrt{q_{1}^{2}+q_{2}^{2}+q_{3}^{2}+q_{4}^{2}},$
generalise the conjugate and modulus of a complex number $x+iy$, and allow the inner product (and associated norm) to be extended to $\mathbb{H}$ as follows. We note that
$\overline{ab}=\overline{b}\,\overline{a},\qquad a,b\in\mathbb{H}\quad\Longrightarrow\quad(AB)^{*}=B^{*}A^{*}\quad\hbox{(for matrices over $\mathbb{H}$)}.$
Definition 1.1
Let $\cal V$} be a finite-dimensional (right) vector space over $\mathbb{F}=\mathbb{R},\mathbb{C},\mathbb{H}$. Then an $\mathbb{F}$-valued map $\langle\cdot,\cdot\rangle:{\cal V}\times{\cal V}\to\mathbb{F}$ is called an inner product if it satisfies
  1. 1.
    Conjugate symmetry: $\langle v,w\rangle=\overline{\langle w,v\rangle}$.
  2. 2.
    Linearity in the second variable: $\langle v,w\beta\rangle=\langle v,w\rangle\beta$, $\langle u,v+w\rangle=\langle u,v\rangle+\langle u,w\rangle$.
  3. 3.
    Positive definiteness: $\langle v,v\rangle>0$, $v\neq 0$.
for all vectors $v,w,u\in{\cal V}$ and scalars $\beta\in\mathbb{F}$.
We will say that $\cal V$} is a real, complex or quaternionic inner product space (respectively). The theory of inner product spaces evolves as in the real and complex cases, though it is not well known, e.g., the Cauchy-Schwarz inequality
$|\langle v,w\rangle|\leq\|v\|\|w\|,\qquad\|v\|:=\sqrt{\langle v,v\rangle},$ (1.1)
holds (with equality if and only if $v$ and $w$ are linearly dependent), though it is not mentioned in the monograph [Rod14]. I think this is in part due to the fact that real and complex valued inner products are often also defined on $\mathbb{H}$-vector spaces. A good treatment is given in [Coh80] (“unitary inner products”) and [GMP13] (“Hermitian quaternionic scalar products”, which includes Cauchy-Schwarz). The prototype of such an inner product is the Euclidean (or standard) inner product
$\langle v,w\rangle:=v^{*}w=\sum_{j}\overline{v_{j}}w_{j},\qquad v,w\in\mathbb{H}^{d}.$ (1.2)
Throughout, we will use the notation $\langle v,w\rangle$ for the Euclidean inner product, sometimes writing $\langle v,w\rangle_{\mathbb{F}}$ to emphasize when all the entries of vectors $v$ and $w$ are in $\mathbb{F}=\mathbb{R},\mathbb{C},\mathbb{H}$. The Euclidean inner product on the entries of a matrix is the Frobenius inner product
$\langle A,B\rangle_{F}:=\mathop{\rm trace}\nolimits(A^{*}B),\qquad\|A\|_{F}^{2}=\langle A,A\rangle_{F}=\sum_{j}\sum_{k}|a_{jk}|^{2}.$ (1.3)
In light of the noncommutativity of the quaternions, we note that scalars come outside an inner product (as we have defined it) as follows
$\langle v\alpha,w\beta\rangle=\overline{\alpha}\langle v,w\rangle\beta.$ (1.4)
The notion of orthogonality, and the Gram-Schmidt process extends in the obvious fashion. There is no need for notions of “left” and “right” orthogonality, since
$\langle v,w\rangle=0\quad\Longleftrightarrow\quad\langle w,v\rangle=\overline{\langle v,w\rangle}=0.$
The Riesz representation also extends to inner products over $\mathbb{H}$, and so the adjoint of a linear map $T:{\cal V}\to{\cal W}$ between finite-dimensional inner product spaces can be defined as the unique linear map $T^{*}:{\cal W}\to{\cal V}$ satisfying
$\langle T^{*}w,v\rangle=\langle w,Tv\rangle,\qquad\forall v\in{\cal V},\ w\in{\cal W}.$
If $T$ and $T^{*}$ are represented as matrices $[T]$ and $T^{*}]$ with respect to orthonormal bases $(v_{j})$ and $(w_{k})$, so that $v=\sum_{j}v_{j}\langle v_{j},v\rangle$, $\forall v\in{\cal V}$, and $w=\sum_{k}w_{k}\langle w_{k},w\rangle$, $\forall w\in{\cal W}$, then
$T]_{jk}=\langle w_{j},Tv_{k}\rangle=\langle T^{*}w_{j},v_{k}\rangle=\overline{\langle v_{k},T^{*}w_{j}\rangle}=\overline{[T^{*}]_{kj}},$
and hence the matrix $T^{*}]$ is the conjugate transpose of the matrix $[T]$. For this reason, it is often assumed that the inner product is the standard inner product on $\mathbb{H}^{d}$, and all calculations are done with matrices, with $A^{*}$ defined to be the conjugate transpose (or Hermitian transpose) of the matrix $A$, as is the case in [Rod14]. The adjoint and Hermitian transpose satisfy some (but not all) of the usual properties, including
$(AB)^{*}=B^{*}A^{*},\qquad(A+B)^{*}=A^{*}+B^{*},\qquad(A^{*})^{*}=A,$
$(A^{*})^{-1}=(A^{-1})^{*}\quad\hbox{(for $A$ invertible)}.$
Note that the transpose does not satisfy $(AB)^{T}=B^{T}A^{T}$ (since $\mathbb{H}$ is not commutative). It can be shown that if $AB=I$ for square matrices over $\mathbb{H}$ then $BA=I$, and so a right inverse exists for $A$ if and only if a left inverse exists, and these inverses are equal (and denoted by $A^{-1}$).
One subtle point, which is not obvious from the matrix formulation, is that scalar multiplication by $\beta\in\mathbb{H}\setminus\mathbb{R}$, i.e., $R_{\beta}:{\cal V}\to{\cal V}:v\mapsto v\beta$ is not an $\mathbb{H}$-linear map, since
$R_{\beta}(v\alpha)=(v\alpha)\beta=v(\alpha\beta)\neq v(\beta\alpha)=(v\beta)\alpha=(R_{\beta}v)\alpha\qquad\hbox{(in general)}.$
Left multiplication of $\mathbb{H}^{d}$ by $\beta$ defines an $\mathbb{H}$-linear map $L_{\beta}:\mathbb{H}^{d}\to\mathbb{H}^{d}:v\mapsto\beta v$, but this is dependent on a choice of basis: it is the linear map which maps $e_{j}\mapsto e_{j}\beta$, i.e., the linear map whose matrix representation with respect to the standard basis $(e_{j})$ is $\beta I$ (see the discussion of [GMP13] §3.1). On the other hand, multiplication of a fixed vector $v\in{\cal V}$ by scalars, i.e., $v]:\mathbb{H}\to{\cal V}:\alpha\mapsto v\alpha$ is an $\mathbb{H}$-linear map:
$v](\beta_{1}\alpha_{1}+\beta_{2}\alpha_{2})=v(\beta_{1}\alpha_{1}+\beta_{2}\alpha_{2})=(v\beta_{1})\alpha_{1}+(v\beta_{2})\alpha_{2}=([v]\beta_{1})\alpha_{1}+([v]\beta_{2})\alpha_{2}.$
Its adjoint $v]^{*}:{\cal V}\to\mathbb{H}$ is given by $v]^{*}=\langle v,\cdot\rangle$, since
$\langle\alpha,[v]^{*}w\rangle=\langle[v]\alpha,w\rangle=\langle v\alpha,w\rangle=\overline{\alpha}\langle v,w\rangle=\overline{\alpha}\langle 1,\langle v,w\rangle\rangle=\langle\alpha,\langle v,w\rangle\rangle.$
The map $[v]$ is sometimes abbreviated simply as $v$, especially when $v\in\mathbb{H}^{d}$ is thought of as a column vector, i.e., as an element of $\mathbb{H}^{d\times 1}$. More generally, a synthesis map
$V=[v_{1},\ldots,v_{n}]:\mathbb{H}^{n}\to{\cal V}:a\mapsto v_{1}a_{1}+\cdots+v_{n}a_{n},$
for a sequence of vectors $v_{1},\ldots,v_{n}\in{\cal V}$, has adjoint the analysis map
$V^{*}:{\cal V}\to\mathbb{H}^{n}:v\mapsto(\langle v_{j},v\rangle)_{j=1}^{n}.$
2 Tight frames
[…]
$A\|v\|^{2}\leq\sum_{j}|\langle v_{j},v\rangle|^{2}\leq B\|v\|^{2},\qquad\forall v\in{\cal H},$ (2.5)
where $A,B>0$ are constants. From this, a “frame expansion” follows, which takes a particularly simple form when $A=B$, i.e.,
$v={1\over A}\sum_{j}v_{j}\langle v_{j},v\rangle,\qquad\forall v\in{\cal H}.$
A prominent early example of the use of such “generalised orthonormal bases” is in the theory of wavelets. Recently, frames have been considered for quaternionic Hilbert space, see, e.g., [KTS17], [VSSS20] (which deal primarily with the frame operator and the construction of dual frames), [KPS21] (equiangular lines and Hadamard matrices), and [NS22] which considers the connectedness of the algebraic variety of quaternionic frames with given norms. Here we consider tight frames (where the dual frame is the frame itself) with a particular emphasis on the classification and construction of such frames. This is related to earlier work of Hoggar [Hog77], [Hog82] and others, which implicitly considers tight frames over quaternionic (and even octonionic) Hilbert spaces.
2.1 Tight frames defined and unitary equivalence
[…]
The monograph [Wal18] is a good reference for those parts of the theory of finite tight frames which we now develop. First we consider equivalent conditions for being a tight frame. For this, we need the polarisation identity for quaternionic Hilbert space. Since this is not well known, we provide it with proof.
Lemma 2.1
(Polarisation identity) For an inner product space over $\mathbb{F}=\mathbb{R},\mathbb{C},\mathbb{H}$, we have
$\langle v,w\rangle={1\over 4}\sum_{r=0}^{m-1}\Bigl{(}\|vi_{r}+w\|^{2}-\|vi_{r}-w\|^{2}\Bigr{)}i_{r},$
where $m=\dim_{\mathbb{R}}(\mathbb{F})$, $(i_{0},i_{1},i_{2},i_{3})=(1,i,j,k)$, and $\langle\cdot,\cdot\rangle$ is linear in the second variable.
Proof: We first observe that for a quaternion $q=q_{0}+q_{1}i_{1}+q_{2}i_{2}+q_{3}i_{3}$, $q_{r}\in\mathbb{R}$, a calculation gives
$\overline{i_{r}}q+\overline{q}i_{r}=2q_{r},\qquad r=0,1,2,3,$ (2.6)
and we write $(q)_{r}=q_{r}$. Expanding, using the properties of the inner product, gives
$\|vi_{r}\pm w\|^{2}$ $=\langle vi_{r},vi_{r}\rangle+\langle\pm w,\pm w\rangle+\langle vi_{r},\pm w\rangle+\langle\pm w,vi_{r}\rangle$
$=\|v\|^{2}+\|w\|^{2}\pm\overline{i_{r}}\langle v,w\rangle\pm\langle w,v\rangle i_{r},$
so that
$\|vi_{r}+w\|^{2}-\|vi_{r}-w\|^{2}=2\bigl{(}\overline{i_{r}}\langle v,w\rangle+\overline{\langle v,w\rangle}i_{r}\bigr{)}=4(\langle v,w\rangle)_{r},$
which gives the result.
This could also be proved by rewriting equation (3.5.1) of [Rod14] for $A=I$ and $(q_{1},q_{2},q_{3})=(i,j,k)$.
The basic characterisations of normalised tight frames generalise.
Proposition 2.1
[…]
  1. (i)
[…]
    $\|v\|^{2}=\sum_{j}|\langle v_{j},v\rangle|^{2},\qquad\forall v\in{\cal H}.$
  2. (ii)
    The frame operator $S=VV^{*}=I$, i.e., we have the frame expansion
    $v=\sum_{j}v_{j}\langle v_{j},v\rangle,\qquad\forall v\in{\cal H}.$
  3. (iii)
    The Plancherel identity
    $\langle v,w\rangle=\sum_{j}\langle v,v_{j}\rangle\langle v_{j},w\rangle,\qquad\forall v,w\in{\cal H}.$
  4. (iv)
    The Gramian $P=V^{*}V$ is a rank $d$ orthogonal projection, i.e., $^{2}=P, P=P$$^{*}=P$.
[…]
$4(\langle v,w\rangle)_{r}$ $=\|vi_{r}+w\|^{2}-\|vi_{r}-w\|^{2}=\sum_{j}\Bigl{(}|\langle v_{j},vi_{r}+w\rangle|^{2}-|\langle v_{j},vi_{r}-w\rangle|^{2}\Bigr{)}$
$=\sum_{j}\Bigl{(}2\overline{i_{r}}\langle v,v_{j}\rangle\langle v_{j},w\rangle+2\langle w,v_{j}\rangle\langle v_{j},v\rangle i_{r}\Bigr{)}=4\sum_{j}(\langle v,v_{j}\rangle\langle v_{j},w\rangle)_{r}.$
Thus (by the Riesz representation, or since the orthogonal complement of $\mathbb{H}^{d}$ is $\{0\}$)
$\langle v,w\rangle=\sum_{j}\overline{\langle v_{j},v\rangle}\langle v_{j},w\rangle=\langle\sum_{j}v_{j}\langle v_{j},v\rangle,w\rangle\quad\Longrightarrow\quad v=\sum_{j}v_{j}\langle v_{j},v\rangle,$
which is (ii).
[…]
$\langle v_{j},v_{k}\rangle=P_{jk}=\sum_{\ell}P_{j\ell}P_{\ell k}=\sum_{\ell}\langle v_{j},v_{\ell}\rangle\langle v_{\ell},v_{k}\rangle,$
which is the Plancherel identity for $v=v_{j}$ and $w=v_{k}$. The implications then follow by extending the Plancherel identity (using linearity and symmetry of the inner product), and calculating $\mathop{\rm rank}\nolimits(P)=\mathop{\rm trace}\nolimits(P)=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(VV^{*}))=d$, by (ii).
[…]
Proposition 2.2
An $n\times n$ matrix $P$ is the Gramian matrix of a normalised tight frame $V=[v_{1},\ldots,v_{n}]$ for $\mathbb{H}^{d}$ if and only if it is an orthogonal projection matrix of rank $d$.
Proof: We have already seen that a normalised tight frame is determined by its Gramian, which is an orthogonal projection of rank $d$ (Proposition 2.1). It remains only to show that such a matrix $P$ corresponds to a normalised tight frame. Let $v_{j}=Pe_{j}$. Then with the Euclidean norm on $\mathbb{H}^{n}$, we have that
$\langle v_{j},v_{k}\rangle=\langle Pe_{j},Pe_{k}\rangle=\langle e_{j},Pe_{k}\rangle=P_{jk},$
so that $(v_{j})$ is such a tight frame (for its $d$-dimensional span).
A finite sequence of unit vectors $(v_{j})$ (or the lines they represent) are said to be equiangular if
$|\langle v_{j},v_{k}\rangle|^{2}=\lambda=c^{2}=(\cos\theta)^{2},\qquad\forall j\neq k.$ (2.7)
The constants $\lambda$, $c$ and $\theta$ all occur in the literature, and are called the (common) angle.
Example 2.1
Four equiangular lines in $\mathbb{H}^{2}$ with $\lambda={1\over 3}$ are given in [Hog77], namely
$w_{1}$ $={1\over\sqrt{2}}\begin{pmatrix}1\cr j\end{pmatrix},\quad w_{2}={1\over\sqrt{6}}\begin{pmatrix}1-\sqrt{2}i\cr j-\sqrt{2}k\end{pmatrix}$,
$w_{3}$ $={1\over 2\sqrt{3}}\begin{pmatrix}\sqrt{2}+\sqrt{3}+i\cr\sqrt{2}j-\sqrt{3}j+k\end{pmatrix},\quad w_{4}={1\over 2\sqrt{3}}\begin{pmatrix}\sqrt{2}-\sqrt{3}+i\cr\sqrt{2}j+\sqrt{3}j+k\end{pmatrix}$.
The Gramian of these vectors (which are a tight frame for $\mathbb{H}^{2}$) has only complex entries, and so they are unitarily equivalent to an equiangular tight frame for $\mathbb{C}^{2}$. They have the same Gramian as the Weyl-Heisenberg SIC $v_{1}=v$, $v_{2}=Sv$, $v_{3}=\Omega v$, $v_{4}=iS\Omega v$, where
$v={1\over\sqrt{6}}\begin{pmatrix}\sqrt{3+\sqrt{3}}\cr{1\over\sqrt{2}}(1+i)\sqrt{3-\sqrt{3}}\end{pmatrix},\quad S=\begin{pmatrix}0&1\cr 1&0\end{pmatrix},\ \Omega=\begin{pmatrix}1&0\cr 0&-1\end{pmatrix}$.
Therefore, there is a unitary map $U$ with $v_{j}=Uw_{j}$, which we calculate as
$U=\begin{pmatrix}z_{1}&-jz_{1}\cr z_{2}&-kz_{2}\end{pmatrix},\quad z_{1}:={\sqrt{3+\sqrt{3}}\over 2\sqrt{3}}+{\sqrt{3-\sqrt{3}}\over 2\sqrt{3}}i,\quad z_{2}:={\sqrt{3+\sqrt{6}}\over 2\sqrt{3}}-{\sqrt{3-\sqrt{6}}\over 2\sqrt{3}}i.$
Though this first example of quaternionic equiangular lines are not “quaternionic”, we will see that such lines do exist, and they are very intriguing.
2.2 The variational characterisation of tight frames
We now seek to extend the variational characterisation for tight frames [BF03], [Wal03]. For $\mathbb{C}^{d}$, this is most easily and transparently proved from the spectral decomposition of the frame operator using $\mathop{\rm trace}\nolimits(AB)=\mathop{\rm trace}\nolimits(BA)$ (see [Wal18], Theorem 6.1). This trace formula no longer holds over the quaternions, even for $1\times 1$ matrices. Instead, we will use the fact
$\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(AB))=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(BA)),$ (2.8)
which follows from the special case $\mathop{\rm Re}\nolimits(ab)=\mathop{\rm Re}\nolimits(ba)$, $\forall a,b\in\mathbb{H}$.
The general spectral theory of matrices over $\mathbb{H}$ is fraught (see [Rod14]), since
$Av=v\lambda\quad\Longrightarrow\quad A(v\alpha)=(v\alpha)\alpha^{-1}\lambda\alpha,$
so that if $v$ is a (right) eigenvector for $\lambda$, then $v\alpha$ is an eigenvector for eigenvalue $\alpha^{-1}\lambda\alpha$. However, Hermitian matrices (those with $A^{*}=A$), have real eigenvalues and are unitarily diagonalisable, as in the complex case.
Lemma 2.2
Let $V=[v_{1},\ldots,v_{n}]$ be vectors in $\mathbb{F}^{d}$, with frame operator $S=VV^{*}$ and Gramian $G=V^{*}V$. Then $\mathop{\rm trace}\nolimits(S^{k})=\mathop{\rm trace}\nolimits(G^{k})$, $k=1,2,\ldots$. In particular,
$\mathop{\rm trace}\nolimits(S)=\sum_{j}\|v_{j}\|^{2},\qquad\mathop{\rm trace}\nolimits(S^{2})=\sum_{j}\sum_{k}|\langle v_{j},v_{k}\rangle|^{2}.$ (2.9)
Proof: The trace of an Hermitian matrix $A$ is real, since $\overline{\langle Ax,x\rangle}=\langle x,Ax\rangle=\langle Ax,x\rangle$. Since $S^{k}$ and $G^{k}$ are Hermitian, they have real trace, and so by (2.8), we have
$\mathop{\rm trace}\nolimits(S^{k})$ $=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(VV^{*}(VV^{*})^{k-1}))=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(V^{*}(VV^{*})^{k-1}V))$
$=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits((V^{*}V)^{k}))=\mathop{\rm trace}\nolimits(G^{k}).$
The formulas for $\mathop{\rm trace}\nolimits(G)$ and $\mathop{\rm trace}\nolimits(G^{2})$ given on the left hand side of (2.9) are easily calculated from $(G)_{jk}=\langle v_{j},v_{k}\rangle$.
Theorem 2.1
(Variational characterisation) Let $v_{1},\ldots,v_{n}$ be vectors in $\mathbb{F}^{d}$, which are not all zero. Then
$\sum_{j=1}^{n}\sum_{k=1}^{n}|\langle v_{j},v_{k}\rangle|^{2}\geq{1\over d}\Bigl{(}\sum_{j=1}^{n}\|v_{j}\|^{2}\Bigr{)}^{2},$ (2.10)
with equality if and only if $(v_{j})_{j=1}^{n}$ is a tight frame for $\mathbb{F}^{d}$.
Proof: Let $V=[v_{j}]$. Since $S=VV^{*}$ is positive definite, it is unitarily diagonalisable $S=U\Lambda U^{*}$, $\Lambda=\mathop{\rm diag}\nolimits(\lambda_{j})$, with real eigenvalues $\lambda_{1},\ldots,\lambda_{d}\geq 0$. From (2.8), we have
$\mathop{\rm trace}\nolimits(S^{k})=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(U\Lambda^{k}U^{*}))=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(\Lambda^{k}U^{*}U))=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(\Lambda^{k}))=\mathop{\rm trace}\nolimits(\Lambda^{k}).$
Thus, the Cauchy-Schwarz inequality gives
$\mathop{\rm trace}\nolimits(S)^{2}=(\sum_{j}\lambda_{j})^{2}=\langle(1),(\lambda_{j})\rangle^{2}\leq\|(1)\|^{2}\|(\lambda_{j})\|^{2}=d\sum_{j}\lambda_{j}^{2}=d\mathop{\rm trace}\nolimits(S^{2}),$
which, by (2.9), is (2.10), with equality if and only if $\lambda_{j}=A$, $\forall j$, $A>0$, i.e.,
$S=U(AI)U^{*}=AI\quad\Longleftrightarrow\quad\hbox{$(v_{j})$ is a tight frame for $\mathbb{F}^{d}$.}$
Note above, since one vector is nonzero, $S=\sum_{j}v_{j}v_{j}^{*}\neq 0$, and so $A\neq 0$.
This variational characterisation of tight frames depends only on the Gramian, and hence the frame up to unitary equivalence. It is easy to verify, and plays a key role in Theorems 3.1 and 3.2. We now consider its implications for equiangular lines.
2.3 Bounds on equiangular lines
We recall that unit vectors $(v_{j})$ in $\mathbb{F}^{d}$ are equiangular if they satisfy (2.7), i.e.,
$|\langle v_{j},v_{k}\rangle|^{2}=\lambda=c^{2}=(\cos\theta)^{2},\qquad\forall j\neq k.$
Those of the most interest have the maximum separation of the corresponding lines, i.e., $\lambda=c^{2}$ small, or, equivalently, $0\leq\theta\leq{\pi\over 2}$ large. Examples that exist in every dimension $d$ are orthonormal bases of $n=d$ vectors ($\lambda=0$, $\theta=90^{\circ}$) and the $n=d+1$ vertices of a regular simplex ($\lambda={1\over d^{2}}$). The formula for the chordal distance
$\rho(v_{j},v_{k}):=\sqrt{1-|\langle v_{j},v_{k}\rangle|^{2}},$
gives a metric on the lines in $\mathbb{H}^{d}$, and accordingly, [CKM16] calls sets of (tight) equiangular lines “(tight) simplices in projective space” (points an equal distance from each other).
As an example of Theorem 2.1, we have the following bound.
Example 2.2
If all the $n$ vectors $(v_{j})$ in $\mathbb{F}^{d}$ have unit norm, then (2.10) reduces to
$\sum_{j=1}^{n}\sum_{k=1}^{n}|\langle v_{j},v_{k}\rangle|^{2}\geq{1\over d}\Bigl{(}\sum_{j=1}^{n}1^{2}\Bigr{)}^{2}={n^{2}\over d}.$
Moreover, if the $(v_{j})$ are equiangular, then the left hand side is $(n^{2}-n)\lambda+n$, and the inequality rearranges to
$\lambda\geq{n-d\over d(n-1)},$ (2.11)
with equality (and maximum possible separation) when the vectors are a tight frame, and for $\lambda<{1\over d}$ it rearranges to the relative bound for equiangular lines
$n\leq{1-\lambda\over{1\over d}-\lambda},\qquad\lambda<{1\over d}.$
The next bound (which is well known for $\mathbb{F}=\mathbb{R},\mathbb{C}$) depends on the underlying field.
Theorem 2.2
Suppose $d>1$. Let $(v_{j})$ be a sequence of $n$ non-parallel unit vectors in $\mathbb{F}^{d}$ giving a set of $n$ equiangular lines, then the orthogonal projections
$P_{j}=v_{j}v_{j}^{*}:v\mapsto v_{j}\langle v_{j},v\rangle,\qquad j=1,\ldots,n,$
are linearly independent over $\mathbb{R}$, and hence
$n\leq\begin{cases}{1\over 2}d(d+1),&\mathbb{F}=\mathbb{R};\\ d^{2},&\mathbb{F}=\mathbb{C};\\ 2d^{2}-d,&\mathbb{F}=\mathbb{H},\end{cases}$ (2.12)
with equality if and only if $(P_{j})$ is a basis for the $\mathbb{R}$-vector space of Hermitian matrices. In these cases, the angle is
$\lambda=\begin{cases}{1\over d+2},&\mathbb{F}=\mathbb{R};\\ {1\over d+1},&\mathbb{F}=\mathbb{C};\\ {1\over d+{1\over 2}},&\mathbb{F}=\mathbb{H}.\end{cases}$ (2.13)
Proof: Since $d>1$, the equiangularity constant $\lambda$ is less than $1$. Using (2.8), we calculate
$\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}))$ $=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(v_{j}v_{j}^{*}v_{k}v_{k}^{*}))$
$=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(v_{j}^{*}v_{k}v_{k}^{*}v_{j}))=|\langle v_{j},v_{k}\rangle|^{2}=\lambda,\qquad j\neq k.$
The $\mathbb{R}$-linear combination $\sum_{j}c_{j}P_{j}$ is Hermitian, and hence its Frobenius norm satisfies
$\|\sum_{j}c_{j}P_{j}\|_{F}^{2}$ $=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(\sum_{j}c_{j}P_{j}\sum_{k}c_{k}P_{k}))=\sum_{j}\sum_{k}c_{j}{c_{k}}\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}))$
$=\sum_{j}\sum_{k}c_{j}{c_{k}}\lambda+\sum_{j}c_{j}{c_{j}}(1-\lambda)=\lambda\Bigl{(}\sum_{j}c_{j}\Bigr{)}^{2}+(1-\lambda)\sum_{j}c_{j}^{2},$
which is zero only for the trivial linear combination.
The $n$ projections $\{P_{j}\}$ belong to the real vector space of $d\times d$ Hermitian matrices which has dimension given by the right hand side of (2.12). For example, for $\mathbb{F}=\mathbb{H}$ the Hermitian matrices are determined by their real diagonal, and the entries above it which can be any quaternions, giving a dimension of $d+{1\over 2}(d^{2}-d)\cdot 4=2d^{2}-d$.
This result for $\mathbb{H}^{d}$, the inequality (2.12), is given in [Hog76], without proof, and as Proposition 2.2 in [CKM16] (which also includes the octonionic case $\mathbb{O}^{3}$).
We are now in a position to discuss quaternionic equiangular lines. We first observe:
  • Quaternionic equiangular lines do exist (for $\lambda<1$, $d>1$).
You will recall from Example 2.1 that Hoggar’s example of four equiangular lines in $\mathbb{H}^{2}$ were in fact lines in $\mathbb{C}^{2}$ (most likely the very first occurrence of a SIC in the literature). For $d=1$, any sequence of unit quaternions is an equiangular tight frame (with $\lambda=1$), which is quaternionic if any ratio of the quaternions is not a complex number. Even though this is a trivial example, we will be able to use such frames to construct unit-norm tight frames in $\mathbb{C}^{2}$ and $\mathbb{R}^{4}$ (Example 3.5). We now give a simple example in $\mathbb{H}^{2}$.
Example 2.3
(Five equiangular lines in $\mathbb{H}^{2}$). Fix $0<t<1$, and consider the four unit vectors
$v_{r}=\begin{pmatrix}t\cr\sqrt{1-t^{2}}\,i_{r}\end{pmatrix},\qquad i_{1}=1,\quad i_{2}=i,\quad i_{3}=j,\quad i_{4}=k.$
These are equiangular, with
$|\langle v_{r},v_{s}\rangle|^{2}=\lambda:=t^{4}+(1-t^{2})^{2},\qquad j\neq k,$
where ${1\over 2}\leq\lambda<1$. By Theorem 2.2, the maximal number of equiangular lines in $\mathbb{C}^{2}$ is four, with $\lambda={1\over 3}$, so these lines are quaternionic. For the maximal separation $\lambda={1\over 2}$, we may add a fifth equiangular line, to obtain five equiangular lines in $\mathbb{H}^{2}$ given by
${1\over\sqrt{2}}\begin{pmatrix}1\cr 1\end{pmatrix},\quad{1\over\sqrt{2}}\begin{pmatrix}1\cr i\end{pmatrix},\quad{1\over\sqrt{2}}\begin{pmatrix}1\cr j\end{pmatrix},\quad{1\over\sqrt{2}}\begin{pmatrix}1\cr k\end{pmatrix},\quad\begin{pmatrix}1\cr 0\end{pmatrix}\ \bigl{(}\hbox{or }\begin{pmatrix}0\cr 1\end{pmatrix}\bigr{)}$. (2.14)
These lines are not tight, since they do not give equality in (2.11), i.e.,
$\lambda={1\over 2}>{3\over 8}={5-2\over 2(5-1)}={n-d\over d(n-1)}.$
They appear exactly as above in [ET20], for the parameter choice $c={1\over\sqrt{2}}$, $\omega={\pi\over 4}$, $\alpha=0$ and $\gamma={\pi\over 4}$.
Taking the five lines of (2.14) and their orthogonal complement gives five MUBs (mutually unbiased bases) in $\mathbb{H}^{2}$, which is a tight frame of ten vectors (see [BW25]).
Another method to obtain tight equiangular lines is via the complementary tight frame. The construction is as follows. Let $G$ be the Gramian of $n>d$ equiangular unit vectors in $\mathbb{F}^{d}$ at an angle $\lambda={n-d\over d(n-1)}\neq 0$, so that $P={d\over n}G$ is an orthogonal projection matrix (Proposition 2.1). The complementary orthogonal projection Q=IPQ=I-P gives an equiangular tight frame of $n$ vectors for $\mathbb{F}^{n-d}$ with Gramian $G_{c}$ given by
$G_{c}={n\over n-d}I-{d\over n-d}G,$
and common angle $\lambda_{c}={d^{2}\over(n-d)^{2}}\lambda={d\over(n-d)(n-1)}$. The equivalent construction for lines is called the Gale dual in [CKM16] (see Corollary 2.12).
Let $c_{d}$ be the right hand side of (2.12), which we can write as
$c_{d}=d+{1\over 2}d(d-1)\cdot m,\qquad m:=\dim_{\mathbb{R}}(\mathbb{F}).$
Since the complementary tight frame also must satisfy the bound (2.12), for $n-d\neq 1$, we have that an equiangular tight frame of $n>d+1$ unit vectors in $\mathbb{F}^{d}$ must satisfy
$n\leq\min\{c_{d},c_{n-d}\}.$ (2.15)
This gives the following (see Theorem 2.18 of [KF08]).
Proposition 2.3
Let $d\geq 2$. An equiangular tight frame of $n>d+1$ vectors for $\mathbb{F}^{d}$ satisfies
$d+{1\over 2}+{\sqrt{{8\over m}d+1}\over 2}\leq n\leq d+{m\over 2}d(d-1),\qquad m=\dim_{\mathbb{R}}(\mathbb{F}),$ (2.16)
so that
$n\geq d+2+j,\qquad\hbox{for }\quad d>{m\over 2}(j+1)(j+2).$ (2.17)
Proof: The condition $n\leq c_{n-d}$ in (2.15) can be written as
$n^{2}-(2d+1)n+d(d+1)-{2\over m}d\geq 0.$
By considering the roots of this quadratic polynomial in $n$, this is satisfied if and only if
$n\leq d+{1\over 2}-{\sqrt{{8\over m}d+1}\over 2}<d,\quad\hbox{or}\quad n\geq d+{1\over 2}+{\sqrt{{8\over m}d+1}\over 2},$
which gives the lower bound in (2.16). The upper bound is the condition $n\leq c_{d}$.
Rearranging the right hand inequality in
$n\geq d+{1\over 2}+{\sqrt{{8\over m}d+1}\over 2}\geq d+2+j,$
gives
$d\geq{m\over 8}\bigl{(}(2j+3)^{2}-1\bigr{)}={m\over 2}(j+1)(j+2),$
which gives (2.17).
The lower bound in (2.16) is a decreasing function of $m$ and the upper bound is an increasing function of $m$. This says that there is more room in $\mathbb{H}^{d}$ for tight equiangular lines than there is in $\mathbb{C}^{d}$, and in turn $\mathbb{R}^{d}$.
Example 2.4
(Five tight equiangular lines in $\mathbb{H}^{3}$) By Proposition 2.3, there cannot be five tight equiangular lines in $\mathbb{R}^{3}$ or $\mathbb{C}^{3}$, but they could exist in $\mathbb{H}^{3}$. We now construct such lines as a complementary tight frame. The following five tight equiangular lines in $\mathbb{H}^{2}$ with $\lambda={3\over 8}$ are given by [ET20]
$V=\begin{pmatrix}1&{\sqrt{3}\over 2\sqrt{2}}&{\sqrt{3}\over 2\sqrt{2}}&{\sqrt{3}\over 2\sqrt{2}}&{\sqrt{3}\over 2\sqrt{2}}\cr 0&{\sqrt{5}\over 2\sqrt{2}}&-{\sqrt{5}\over 6\sqrt{2}}+{\sqrt{5}\over 3}i&-{\sqrt{5}\over 6\sqrt{2}}-{\sqrt{5}\over 6}i+{\sqrt{5}\over 2\sqrt{3}}j&-{\sqrt{5}\over 6\sqrt{2}}-{\sqrt{5}\over 6}i-{\sqrt{5}\over 2\sqrt{3}}j\end{pmatrix}$.
The complementary tight frame therefore gives five equiangular lines in $\mathbb{H}^{3}$ at angle $\lambda={1\over 6}$. A concrete presentation of these lines is
$W=\begin{pmatrix}1&-{1\over\sqrt{6}}&-{1\over\sqrt{6}}&-{1\over\sqrt{6}}&-{1\over\sqrt{6}}\cr 0&{\sqrt{5}\over\sqrt{6}}&-{\sqrt{5}\over 3\sqrt{6}}-{\sqrt{5}\over 3\sqrt{3}}i&-{\sqrt{5}\over 3\sqrt{6}}+{\sqrt{5}\over 6\sqrt{3}}i-{\sqrt{5}\over 6}j&-{\sqrt{5}\over 3\sqrt{6}}+{\sqrt{5}\over 6\sqrt{3}}i+{\sqrt{5}\over 6}j\cr 0&0&{\sqrt{5}\over 3}&-{\sqrt{5}\over 6}+{\sqrt{5}\over 2\sqrt{3}}k&-{\sqrt{5}\over 6}-{\sqrt{5}\over 2\sqrt{3}}k\end{pmatrix}$.
This was obtained by the following general method. The condition $VV^{*}=AI$ for $V$ to be a tight frame is that the entrywise conjugates of the rows of $V$ are orthogonal and of equal length, i.e., $V^{*}$ has orthogonal columns of equal length. By using Gram-Schmidt, add orthogonal columns of equal length to obtain $V^{*},W^{*}]$ a scalar multiple of a unitary matrix. Then $W$ is a tight frame, which is the complement of $V$, since
$\begin{pmatrix}V^{*}&W^{*}\end{pmatrix}\begin{pmatrix}V^{*}&W^{*}\end{pmatrix}^{*}=\begin{pmatrix}V^{*}&W^{*}\end{pmatrix}\begin{pmatrix}V\cr W\end{pmatrix}=V^{*}V+W^{*}W=AI.$
Above we used the fact that the columns of the square matrix $\begin{pmatrix}V^{*}&W^{*}\end{pmatrix}$ over $\mathbb{H}$ are orthogonal. For frames over $\mathbb{C}$, this is equivalent to the rows being orthogonal. For frames over the quaternions, it is necessary to make this distinction. Indeed, there exist unitary matrices (orthogonal columns) whose rows are not orthogonal, e.g.,
$U:={1\over\sqrt{2}}\begin{pmatrix}1&i\cr j&k\end{pmatrix},\qquad U^{*}U=UU^{*}=\begin{pmatrix}1&0\cr 0&1\end{pmatrix},\qquad(U^{T})^{*}(U^{T})=\begin{pmatrix}1&j\cr-j&1\end{pmatrix}$.
Example 2.5
By Proposition 2.3, there cannot be six tight equiangular lines in $\mathbb{R}^{4}$ or $\mathbb{C}^{4}$, but they do exist in $\mathbb{H}^{4}$, by taking the complementary tight frame to the six tight equiangular lines in $\mathbb{H}^{2}$ of [KF08], [ET20] (obtained independently).
We now consider tight equiangular lines in general, before giving a striking summary of the known results for two dimensions. For $n$ tight equiangular lines in $\mathbb{H}^{d}$ (or $\mathbb{C}^{d}$, $\mathbb{R}^{d}$), the angle is
$\lambda={n-d\over d(n-1)},\qquad n>d,$
with the following specific cases (in order of the number of vectors)
$\lambda=0\quad\hbox{(orthonormal basis)},\qquad\lambda={1\over d^{2}}\quad\hbox{(vertices of a simplex)},$
and sets of lines giving the bounds of Theorem 2.2
$\lambda={1\over d+2},\qquad\lambda={1\over d+1}\quad\hbox{(SIC)},\qquad\lambda={1\over d+{1\over 2}}\quad\hbox{(maximal set of lines in $\mathbb{H}^{d}$}).$
The theory as is stands does not preclude the bounds above being reached by lines from a larger space, e.g., $n={1\over 2}d(d+1)$ complex or even quaternionic lines in $\mathbb{H}^{d}$. This does not occur for two dimensions. Since
$\Bigl{(}{\partial\lambda\over\partial n}\Bigr{)}_{d}={d-1\over d(n-1)^{2}}>0,$
$\lambda$ increases with the number of tight equiangular lines $n$ (for $d$ fixed), taking the possible values
$\lambda=0,{1\over d^{2}},\ldots,{1\over d+2},\ldots,{1\over d+1},\ldots,{2\over 2d+1}.$
Equiangular lines are classified up to projective unitary equivalence (see Section 5).
In two dimensions, the tight equiangular lines given by an orthonormal basis, the Mercedes-Benz frame and the SIC (two, three and four vectors, respectively) are well known, as is their uniqueness in $\mathbb{C}^{2}$. Putting these examples together with the five and six sets of equianglar lines of [KF08], [ET20] gives a complete characterisation of equiangular lines in $\mathbb{H}^{2}$.
Theorem 2.3
There is a unique set of $n$ tight equiangular lines in $\mathbb{H}^{2}$ for $n=2,3,4,5,6$, with corresponding angles $\lambda=0,{1\over 4},{1\over 3},{3\over 8},{2\over 5}$.
2.4 Equi-isoclinic and equichordal subspaces
We now consider generalisations of equiangularity of lines to $r$-subspaces ($r$-dimensional subspaces). Let $P_{j}$ and $P_{k}$ be the orthogonal projections onto $r$-subspaces $V_{j}$ and $V_{k}$. Then
$\|P_{j}-P_{k}\|_{F}^{2}=\mathop{\rm trace}\nolimits((P_{j}-P_{k})^{2})=2r-\mathop{\rm trace}\nolimits(P_{j}P_{k}+P_{k}P_{j})\geq 0.$
For $\mathbb{F}^{d}=\mathbb{R}^{d},\mathbb{C}^{d}$, we have $\mathop{\rm trace}\nolimits(P_{j}P_{k})=\mathop{\rm trace}\nolimits(P_{k}P_{j})\in\mathbb{R}$, and a collection of $r$-subspaces is said to be equichordal (see [FJMW17]) if the corresponding orthogonal projections satisfy
$\langle P_{j},P_{k}\rangle_{F}=\mathop{\rm trace}\nolimits(P_{j}P_{k})=\lambda r,\qquad j\neq k,$
which reduces to the equiangularity condition (2.7) in the case of lines ($r=1$).
For $\mathbb{H}^{d}$, $\mathop{\rm trace}\nolimits(P_{j}P_{k})$ need not be real, nor equal to $\mathop{\rm trace}\nolimits(P_{k}P_{j})$, e.g., for
$P={1\over 2}\begin{pmatrix}1&-i\cr i&1\end{pmatrix},\quad Q={1\over 2}\begin{pmatrix}1&-j\cr j&1\end{pmatrix},\qquad PQ={1\over 4}\begin{pmatrix}1-k&-i-j\cr i+j&1-k\end{pmatrix}$,
and $\mathop{\rm trace}\nolimits(PQ)={1\over 2}(1-k)\neq{1\over 2}(1+k)=\mathop{\rm trace}\nolimits(QP)$. However, by (2.8), we do have
$\mathop{\rm trace}\nolimits(P_{j}P_{k}+P_{k}P_{j})=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}+P_{k}P_{j}))=2\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k})),$
and so we say that $r$-subspaces in $\mathbb{H}^{d}$ (or $\mathbb{R}^{d},\mathbb{C}^{d}$) are equichordal if
$\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}))=\lambda r,\quad j\neq k\quad\Longleftrightarrow\quad\|P_{j}-P_{k}\|_{F}^{2}=2(1-\lambda)r,\quad j\neq k.$ (2.18)
Two $r$-subspaces ${\cal V}_{j}$ and ${\cal V}_{k}$, $j\neq k$, are isoclinic with parameter $0\leq\lambda\leq 1$ (see [LS73], [Hog77], [FIJM24]) if the orthogonal projection $P_{jk}$ onto ${\cal V}_{j}+{\cal V}_{k}$ satisfies
$(1-\lambda)P_{jk}=(P_{j}-P_{k})^{2}.$
An equivalent condition to being isoclinic is
$P_{j}P_{k}P_{j}=\lambda P_{j},\qquad P_{k}P_{j}P_{k}=\lambda P_{k},\qquad j\neq k,$ (2.19)
which follows from the observation
$(1-\lambda)P_{j}=(P_{j}-P_{k})^{2}P_{j}\quad\Longleftrightarrow\quad P_{j}P_{k}P_{j}=\lambda P_{j}.$
Hoggar [Hog77] claims that just one of the conditions (2.19) is required (over $\mathbb{H}$), which follows by writing $P_{j}=V_{j}V_{j}^{*}$, $V_{j}^{*}V_{j}=I$, and the implications
$P_{j}P_{k}P_{j}=\lambda P_{j}\quad\Longleftrightarrow\quad(V_{j}^{*}V_{k})(V_{k}^{*}V_{j})=\lambda I\quad\Longleftrightarrow\quad V_{k}^{*}V_{j}V_{j}^{*}V_{k}=\lambda I\quad\Longleftrightarrow\quad P_{k}P_{j}P_{k}=\lambda P_{k}.$
Subspaces $(V_{j})$ are said to be equi-isoclinic with parameter $0\leq\lambda\leq 1$ if (2.19) holds. Equi-isoclinic subspaces are equichordal, since
$P_{j}P_{k}P_{j}=\lambda P_{j}\quad\Longrightarrow\quad\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}))=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}P_{j}))=\mathop{\rm trace}\nolimits(\lambda P_{j})=\lambda r.$
The orthogonal complement $(V_{j}^{\perp})$ of equichordal subspaces is equichordal, since
$\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits((I-P_{j})(I-P_{k})))=d-r-r+\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(P_{j}P_{k}))=d-2r+\lambda r,\quad j\neq k.$
However, the orthogonal complements $(V_{j}^{\perp})$ of equi-isoclinic subspaces $(V_{j})$ are not in general equi-isoclinic, as the following example shows.
Example 2.6
(Two isoclinic planes do not exist in $\mathbb{R}^{3}$). Consider the equi-isoclinic $1$-dimensional subspaces given by
$v_{1}=\begin{pmatrix}1\cr 0\cr 0\end{pmatrix},\quad v_{2}=\begin{pmatrix}\sqrt{1-a^{2}-b^{2}}\cr a\cr b\end{pmatrix}$.
The orthogonal projections $Q_{j}=I-v_{j}v_{j}^{*}$ onto the complementary subspaces satisfy
$Q_{1}Q_{2}Q_{1}=\begin{pmatrix}0&0&0\cr 0&1-a^{2}&-ab\cr 0&-ab&1-b^{2}\end{pmatrix},\qquad Q_{1}=\begin{pmatrix}0&0&0\cr 0&1&0\cr 0&0&1\end{pmatrix}$.
Hence, for $V_{1}^{\perp}$ and $V_{2}^{\perp}$ to be isoclinic, we must have that $a=b=0$, i.e., $V_{1}=V_{2}$. Thus there cannot be two (nonequal) isoclinic planes in $\mathbb{R}^{3}$, despite the fact that there can be up to six equi-isoclinic lines in $\mathbb{R}^{3}$.
3 From $\mathbb{R}$ to $\mathbb{C}$ and $\mathbb{C}$ to $\mathbb{H}$, and back
There is a natural inclusion $\mathbb{R}\subset\mathbb{C}\subset\mathbb{H}$ and hence of $\mathbb{R}^{d}\subset\mathbb{C}^{d}\subset\mathbb{H}^{d}$. Since tight frames are determined up to unitary equivalence by their Gramians:
  • There is a unitary map of a tight frame to $\mathbb{R}^{d}$ if and only if its Gramian has real entries, and we say the tight frame is real.
  • There is a unitary map of a tight frame to $\mathbb{C}^{d}$ if and only if its Gramian has complex entries, and we say the tight frame is complex if its Gramian has a nonreal entry.
  • If the Gramian of a tight frame has a noncomplex entry, then we say that it is a quaternionic tight frame.
As an example, the four equiangular lines in $\mathbb{H}^{2}$ of Hoggar [Hog77] are lines in $\mathbb{C}^{2}$ (see Example 2.1). For tight frames up to projective unitary equivalence, i.e., thought of as lines, the corresponding analogue is more involved, see Section 5.
There is also a natural identification of a point $z=x+iy\in\mathbb{C}$ (in the complex plane) with a point $(x,y)\in\mathbb{R}^{2}$ (in the plane). We generalise this, by defining an invertible $\mathbb{R}$-linear map
$\cdot]_{\mathbb{R}}:\mathbb{C}^{d}\to\mathbb{R}^{2d}:v\mapsto\begin{pmatrix}\mathop{\rm Re}\nolimits v\cr\mathop{\rm Im}\nolimits v\end{pmatrix},\qquad\mathop{\rm Re}\nolimits v={v+\overline{v}\over 2},\quad\mathop{\rm Im}\nolimits v={v-\overline{v}\over 2i}.$ (3.20)
Based on a thorough analysis of this, we will then define an analogous map $\mathbb{H}^{d}\to\mathbb{C}^{2d}$. The first subtle point, is that $\cdot$$_{\mathbb{R}}$ maps $k$-dimensional complex-subspaces of $\mathbb{C}^{d}$ to real $(2k)$-dimensional subspaces of $\mathbb{R}^{2d}$. To see why this is, we first calculate the image of a complex scalar multiple $\alpha+i\beta$ of a vector $v=x+iy$
$(\alpha+i\beta)v=(\alpha+i\beta)(x+iy)=\alpha x-\beta y+i(\alpha y+\beta x),$
which gives
$(\alpha+i\beta)v]_{\mathbb{R}}=\alpha\begin{pmatrix}\mathop{\rm Re}\nolimits v\cr\mathop{\rm Im}\nolimits v\end{pmatrix}+\beta\begin{pmatrix}-\mathop{\rm Im}\nolimits v\cr\mathop{\rm Re}\nolimits v\end{pmatrix}=\alpha[v]_{\mathbb{R}}+\beta[iv]_{\mathbb{R}}.$ (3.21)
Thus the one-dimensional complex subspace spanned by $v\in\mathbb{C}^{d}$ is mapped to the real two-dimensional subspace
$\mathop{\rm span}\nolimits_{\mathbb{C}}\{v\}]_{\mathbb{R}}=\mathop{\rm span}\nolimits_{\mathbb{R}}\{\begin{pmatrix}\mathop{\rm Re}\nolimits v\cr\mathop{\rm Im}\nolimits v\end{pmatrix},\begin{pmatrix}-\mathop{\rm Im}\nolimits v\cr\mathop{\rm Re}\nolimits v\end{pmatrix}\}\qquad\hbox{(orthogonal vectors in $\mathbb{R}^{2d}$)}.$
The general result then follows from the correspondence between linear dependencies
$\sum_{\ell}(\alpha_{\ell}+i\beta_{\ell})v_{\ell}=0\quad\Longleftrightarrow\quad\sum_{\ell}\Bigl{\{}\alpha_{\ell}\begin{pmatrix}\mathop{\rm Re}\nolimits v_{\ell}\cr\mathop{\rm Im}\nolimits v_{\ell}\end{pmatrix}+\beta_{\ell}\begin{pmatrix}-\mathop{\rm Im}\nolimits v_{\ell}\cr\mathop{\rm Re}\nolimits v_{\ell}\end{pmatrix}\Bigr{\}}=0.$
We also calculate
$\langle v,w\rangle$ $=\langle\mathop{\rm Re}\nolimits v+i\mathop{\rm Im}\nolimits v,\mathop{\rm Re}\nolimits w+i\mathop{\rm Im}\nolimits w\rangle$
$=\langle\mathop{\rm Re}\nolimits v,\mathop{\rm Re}\nolimits w\rangle+\langle\mathop{\rm Im}\nolimits v,\mathop{\rm Im}\nolimits w\rangle+i(\langle\mathop{\rm Re}\nolimits v,\mathop{\rm Im}\nolimits w\rangle-\langle\mathop{\rm Im}\nolimits v,\mathop{\rm Re}\nolimits w\rangle),$
so that
$\mathop{\rm Re}\nolimits(\langle v,w\rangle_{\mathbb{C}})=\langle[v]_{\mathbb{R}},[w]_{\mathbb{R}}\rangle_{\mathbb{R}},\quad\mathop{\rm Im}\nolimits(\langle v,w\rangle_{\mathbb{C}})=\langle[iv]_{\mathbb{R}},[w]_{\mathbb{R}}\rangle_{\mathbb{R}},\qquad$ (3.22)
$\langle[v]_{\mathbb{R}},[iv]_{\mathbb{R}}\rangle_{\mathbb{R}}=0.$ (3.23)
Let $A:\mathbb{C}^{n}\to\mathbb{C}^{m}$ a $\mathbb{C}$-linear map be represented as an $\mathbb{R}$-linear map $A]_{\mathbb{R}}:\mathbb{R}^{2n}\to\mathbb{R}^{2m}$ under this identification, i.e., $A]_{\mathbb{R}}:=[\cdot]_{\mathbb{R}}A[\cdot]_{\mathbb{R}}^{-1}$. Then
$A(u+iv)$ $=(\mathop{\rm Re}\nolimits(A)+i\mathop{\rm Im}\nolimits(A))(u+iv)$
$=\mathop{\rm Re}\nolimits(A)u-\mathop{\rm Im}\nolimits(A)v+i\mathop{\rm Im}\nolimits(A)u+i\mathop{\rm Re}\nolimits(A)v,\quad u,v\in\mathbb{R}^{n},$
and $\mathop{\rm Re}\nolimits(A^{*})=\mathop{\rm Re}\nolimits(A)^{T}$, $\mathop{\rm Im}\nolimits(A^{*})=-\mathop{\rm Im}\nolimits(A)^{T}$, so that
$[A]_{\mathbb{R}}$ $=\begin{pmatrix}\mathop{\rm Re}\nolimits(A)&-\mathop{\rm Im}\nolimits(A)\cr\mathop{\rm Im}\nolimits(A)&\mathop{\rm Re}\nolimits(A)\end{pmatrix},\qquad\mathop{\rm rank}\nolimits([A]_{\mathbb{R}})=2\mathop{\rm rank}\nolimits(A),$
$[A^{*}]_{\mathbb{R}}$ $=\begin{pmatrix}\mathop{\rm Re}\nolimits(A)^{T}&\mathop{\rm Im}\nolimits(A)^{T}\cr-\mathop{\rm Im}\nolimits(A)^{T}&\mathop{\rm Re}\nolimits(A)^{T}\end{pmatrix}=[A]_{\mathbb{R}}^{T}.$
The usual rules for matrix multiplication follow, e.g, $A]_{\mathbb{R}}[B]_{\mathbb{R}}=[AB]_{\mathbb{R}}$. One must be careful if a vector $v\in\mathbb{C}^{d}$ is being thought of as a $d\times 1$ matrix, i.e., the linear map $v]:\mathbb{C}\to\mathbb{C}^{d}:\alpha\mapsto\alpha v$, since $v]_{\mathbb{R}}\in\mathbb{R}^{2d\times 1}$, $v]]_{\mathbb{R}}\in\mathbb{R}^{2d\times 2}$. In particular, the familiar formula $P=vv^{*}$ for the orthogonal projection onto a unit vector $v\in\mathbb{C}^{d}$, is $P=[v][v]^{*}$, which maps as follows
$P]_{\mathbb{R}}=[[v]]_{\mathbb{R}}[[v]^{*}]_{\mathbb{R}}=[[v]]_{\mathbb{R}}[[v]]_{\mathbb{R}}^{T},\qquad[[v]]_{\mathbb{R}}=\begin{pmatrix}\mathop{\rm Re}\nolimits v&-\mathop{\rm Im}\nolimits v\cr\mathop{\rm Im}\nolimits v&\mathop{\rm Re}\nolimits v\end{pmatrix}$.
This is the orthogonal projection onto
$\mathop{\rm span}\nolimits_{\mathbb{C}}\{v\}]_{\mathbb{R}}=\mathop{\rm span}\nolimits_{\mathbb{R}}\{[v]_{\mathbb{R}},[iv]_{\mathbb{R}}\},\qquad[v]_{\mathbb{R}}=\begin{pmatrix}\mathop{\rm Re}\nolimits v\cr\mathop{\rm Im}\nolimits v\end{pmatrix},\quad[iv]_{\mathbb{R}}=\begin{pmatrix}-\mathop{\rm Im}\nolimits v\cr\mathop{\rm Re}\nolimits v\end{pmatrix}$.
The identification $\cdot$$_{\mathbb{R}}$ preserves various properties of linear maps, see Theorem 3.3. In particular, orthogonal projections map to orthogonal projections, and hence:
  • Equi-isoclinic subspaces of dimension $r$ in $\mathbb{C}^{d}$ correspond to equi-isoclinic subspaces of dimension $2r$ in $\mathbb{R}^{2d}$, and similarly for equichordal subspaces.
We now consider the situation for tight frames, which is somewhat more involved, e.g., a basis for $\mathbb{C}^{d}$ does not correspond to a basis for $\mathbb{R}^{2d}$ (which has twice the dimension). Let $V=V_{1}+iV_{2}$ be the synthesis map for a sequence of vectors $v_{1},\ldots,v_{n}\in\mathbb{C}^{d}$, and $V_{\mathbb{R}}$ be the corresponding map for the sequence $v_{1}]_{\mathbb{R}},\ldots,[v_{n}]_{\mathbb{R}}\in\mathbb{R}^{2d}$, i.e.,
$V_{\mathbb{R}}=\begin{pmatrix}V_{1}\cr V_{2}\end{pmatrix}\in\mathbb{R}^{2d\times n}.$
Then $V$ gives a tight frame for $\mathbb{C}^{d}$ if and only if
$VV^{*}=(V_{1}+iV_{2})(V_{1}^{*}-iV_{2}^{*})=V_{1}V_{1}^{*}+V_{2}V_{2}^{*}+i(V_{2}V_{1}^{*}-V_{1}V_{2}^{*})=AI,$
where $dA:=\sum_{j}\|v_{j}\|^{2}=\mathop{\rm trace}\nolimits(VV^{*})=\mathop{\rm trace}\nolimits(V_{\mathbb{R}}V_{\mathbb{R}}^{T})=$, i.e.,
$V_{1}V_{1}^{T}+V_{2}V_{2}^{T}=AI,\qquad V_{2}V_{1}^{T}-V_{1}V_{2}^{T}=0,$
and $V_{\mathbb{R}}$ gives a tight frame for $\mathbb{R}^{2d}$ if and only if
$V_{\mathbb{R}}V_{\mathbb{R}}^{*}=\begin{pmatrix}V_{1}\cr V_{2}\end{pmatrix}\begin{pmatrix}V_{1}^{T}&V_{2}^{T}\end{pmatrix}=\begin{pmatrix}V_{1}V_{1}^{T}&V_{1}V_{2}^{T}\cr V_{2}V_{1}^{T}&V_{2}V_{2}^{T}\end{pmatrix}={1\over 2}A\begin{pmatrix}I&0\cr 0&I\end{pmatrix}$,
i.e.,
$V_{1}V_{1}^{T}=V_{2}V_{2}^{T}={1\over 2}AI,\qquad V_{1}V_{2}^{T}=V_{2}V_{1}^{T}=0.$ (3.24)
Thus all tight frames for $\mathbb{R}^{2d}$ map to tight frames for $\mathbb{C}^{d}$, and a tight frame for $\mathbb{C}^{d}$ gives a tight frame for $\mathbb{R}^{2d}$ if and only if (3.24) holds. This condition says that $V_{1}$ and $V_{2}$ are tight frames for $\mathbb{R}^{d}$ (with the same frame bound) which are orthogonal (see [Wal18] §3.5). We now show that (3.24) depends only on $V$ up to unitary equivalence.
Let $U=U_{1}+iU_{2}$ be unitary, then $UU^{*}=U_{1}U_{1}^{T}+U_{2}U_{2}^{T}+i(U_{2}U_{1}^{T}-U_{1}U_{2}^{T})=I$, which is equivalent to
$U_{1}U_{1}^{T}+U_{2}U_{2}^{T}=I,\qquad U_{2}U_{1}^{T}-U_{1}U_{2}^{T}=0.$ (3.25)
Suppose that $V$ satisfies (3.24), then
$UV=[Uv_{1},\ldots,Uv_{n}]=(U_{1}+iU_{2})(V_{1}+iV_{2})=(U_{1}V_{1}-U_{2}V_{2})+i(U_{2}V_{1}+U_{1}V_{2}),$
$A=\sum_{j}\|v_{j}\|^{2}=\sum_{j}\|Uv_{j}\|^{2}$, and using (3.25), we calculate
$\mathop{\rm Re}\nolimits(UV)\mathop{\rm Re}\nolimits(UV)^{T}$ $=(U_{1}V_{1}-U_{2}V_{2})(V_{1}^{T}U_{1}^{T}-V_{2}^{T}U_{2}^{T})={1\over 2}A(U_{1}U_{1}^{T}+U_{2}U_{2}^{T})={1\over 2}AI,$
$\mathop{\rm Im}\nolimits(UV)\mathop{\rm Im}\nolimits(UV)^{T}$ $=(U_{2}V_{1}+U_{1}V_{2})(V_{1}^{T}U_{2}^{T}+V_{2}^{T}U_{1}^{T})={1\over 2}A(U_{2}U_{2}^{T}+U_{1}U_{1}^{T})={1\over 2}AI,$
$\mathop{\rm Re}\nolimits(UV)\mathop{\rm Im}\nolimits(UV)^{T}$ $=(U_{1}V_{1}-U_{2}V_{2})(V_{1}^{T}U_{2}^{T}+V_{2}^{T}U_{1}^{T})={1\over 2}A(U_{1}U_{2}^{T}-U_{2}U_{1}^{T})=0,$
so that $UV$ satisfies (3.24).
Since the condition for a tight frame for $\mathbb{C}^{d}$ to be a tight frame for $\mathbb{R}^{2d}$ depends only on $V$ up to unitary equivalence, it follows that this condition can be written in terms of the Gramian of $V$. The Gramians of $V$ and $V_{\mathbb{R}}$ are
$V^{*}V=(V_{1}^{*}-iV_{2}^{*})(V_{1}+iV_{2})=V_{1}^{T}V_{1}+V_{2}^{T}V_{2}+i(V_{1}^{T}V_{2}-V_{2}^{T}V_{1}),$
$V_{\mathbb{R}}^{*}V_{\mathbb{R}}=\begin{pmatrix}V_{1}^{T}&V_{2}^{T}\end{pmatrix}\begin{pmatrix}V_{1}\cr V_{2}\end{pmatrix}=V_{1}^{T}V_{1}+V_{2}^{T}V_{2}.$
The variational characterisation for being a tight frame for $\mathbb{C}^{d}$ and for $\mathbb{R}^{2d}$ are
$\|V^{*}V\|_{F}^{2}={1\over d}(\mathop{\rm trace}\nolimits(V^{*}V))^{2},\qquad\|V_{\mathbb{R}}^{*}V_{\mathbb{R}}\|_{F}^{2}={1\over 2d}(\mathop{\rm trace}\nolimits(V_{\mathbb{R}}^{*}V_{\mathbb{R}}))^{2}.$
Since $\mathop{\rm trace}\nolimits(V^{*}V)=\mathop{\rm trace}\nolimits(V_{1}^{T}V_{1}+V_{2}^{T}V_{2})=\mathop{\rm trace}\nolimits(V_{\mathbb{R}}^{*}V_{\mathbb{R}})$, a tight frame for $\mathbb{C}^{d}$ gives a tight frame for $\mathbb{R}^{2d}$ if and only if
$2\|V_{\mathbb{R}}^{*}V_{\mathbb{R}}\|_{F}^{2}-\|V^{*}V\|_{F}^{2}=0.$ (3.26)
By writing this explicitly in terms of $V^{*}V$ (cf [Wal20b]), we obtain the following.
Theorem 3.1
Let $\cdot]_{\mathbb{R}}:\mathbb{C}^{d}\to\mathbb{R}^{2d}$ be the correspondence (3.20) between $\mathbb{C}^{d}$ and $\mathbb{R}^{2d}$. Then
  1. 1.
    Tight frames for $\mathbb{R}^{2d}$ correspond to tight frames for $\mathbb{C}^{d}$.
  2. 2.
    A tight frame $V=[v_{1},\ldots,v_{n}]$ for $\mathbb{C}^{d}$ corresponds to a tight frame for $\mathbb{R}^{2d}$ if and only if it satisfies
    $\sum_{j}\sum_{k}\langle v_{j},v_{k}\rangle^{2}=0,$ (3.27)
    which can also be written as
    $\sum_{j}\sum_{k}(\mathop{\rm Re}\nolimits\langle v_{j},v_{k}\rangle)^{2}=\sum_{j}\sum_{k}(\mathop{\rm Im}\nolimits\langle v_{j},v_{k}\rangle)^{2}.$ (3.28)
Proof: In light of our previous discussion, it remains only to show that (3.26) can be written as (3.27) and (3.28). Using (3.22), we have
$\|V_{\mathbb{R}}^{*}V_{\mathbb{R}}\|_{F}^{2}-\|V^{*}V\|_{F}^{2}$ $=2\sum_{j}\sum_{k}\langle[v_{j}]_{\mathbb{R}},[v_{k}]_{\mathbb{R}}\rangle^{2}-\sum_{j}\sum_{k}|\langle v_{j},v_{k}\rangle|^{2}$
$=2\sum_{j}\sum_{k}(\mathop{\rm Re}\nolimits\langle v_{j},v_{k}\rangle)^{2}-\sum_{j}\sum_{k}|\langle v_{j},v_{k}\rangle|^{2}=0.$
By taking $z=\langle v_{j},v_{k}\rangle$ in
$2(\mathop{\rm Re}\nolimits(z))^{2}-|z|^{2}=2\Bigl{(}{z+\overline{z}\over 2}\Bigr{)}^{2}-z\overline{z}={1\over 2}(z^{2}+\overline{z}^{2}),$
we see that this condition can be written as
${1\over 2}\sum_{j}\sum_{k}\bigl{(}\langle v_{j},v_{k}\rangle^{2}+\langle v_{k},v_{j}\rangle^{2}\bigr{)}=\sum_{j}\sum_{k}\langle v_{j},v_{k}\rangle^{2}=0.$
which gives (3.27). By substituting in $|\langle v_{j},v_{k}\rangle|^{2}=(\mathop{\rm Re}\nolimits\langle v_{j},v_{k}\rangle)^{2}+(\mathop{\rm Im}\nolimits\langle v_{j},v_{k}\rangle)^{2},$ we obtain (3.28).
Example 3.1
A tight frame $(z_{j})$ for $\mathbb{C}$ corresponds to a tight frame for $\mathbb{R}^{2}$ if and only if
$\sum_{j}\sum_{k}(z_{j}\overline{z_{k}})^{2}=\Bigl{(}\sum_{j}z_{j}^{2}\Bigr{)}\Bigl{(}\sum_{k}\overline{z_{k}}^{2}\Bigr{)}=\Bigl{|}\sum_{j}z_{j}^{2}\Bigr{|}^{2}=0\quad\Longleftrightarrow\quad\sum_{j}z_{j}^{2}=0.$
The complex number $z_{j}^{2}=(x_{j}+iy_{j})^{2}$ corresponding to a point $(x_{j},y_{j})$ is sometimes called a diagram vector, and the condition that a frame for $\mathbb{R}^{2}$ is tight if and only if its diagram vectors sum to zero is well known.
We now give a map $\mathbb{H}^{d}\to\mathbb{C}^{2d}$ that has similar properties to $\cdot]_{\mathbb{R}}:\mathbb{C}^{d}\to\mathbb{R}^{2d}$. This is based on the following analogue of the polar decomposition for $\mathbb{C}$, the Cayley-Dickson construction, that every quaternion $q\in\mathbb{H}$ can be written uniquely
$q=z+wj,\qquad z,w\in\mathbb{C}.$ (3.29)
Moreover, we observe the “commutativity” relation
$jz=\overline{z}j,\qquad\forall z\in\mathbb{C},$
which implies
$jA=\overline{A}j,\qquad\forall A\in\mathbb{C}^{m\times n}.$
Let $\mathbb{H}^{d}$ be a right vector space, and define a $\mathbb{C}$-linear map
$\cdot]_{\mathbb{C}}:\mathbb{H}^{d}\to\mathbb{C}^{2d}:z+wj\mapsto\begin{pmatrix}z\cr\overline{w}\end{pmatrix}$, (3.30)
The conjugation $\overline{w}$ is necessary for $\mathbb{C}$-linearity: $(z+wj)\alpha=z\alpha+wj\alpha=z\alpha+w\overline{\alpha}j$ gives
$(z+wj)\alpha]_{\mathbb{C}}=\begin{pmatrix}z\alpha\cr\overline{w}\alpha\end{pmatrix}=\begin{pmatrix}z\cr\overline{w}\end{pmatrix}\alpha=[z+wj]_{\mathbb{C}}\alpha\qquad\forall\alpha\in\mathbb{C}.$
Let $\mathop{\rm Co}\nolimits_{1}$ and $\mathop{\rm Co}\nolimits_{2}$ be the $\mathbb{C}$-linear maps $\mathbb{H}^{d}\to\mathbb{C}^{d}$ giving the “complex coordinates” of $q=z+wj$, i.e.,
$\mathop{\rm Co}\nolimits_{1}(z+wj):=z,\qquad\mathop{\rm Co}\nolimits_{2}(z+wj):=\overline{w}.$
We note in particular, that
$|q|^{2}=|\mathop{\rm Co}\nolimits_{1}(q)|^{2}+|\mathop{\rm Co}\nolimits_{2}(q)|^{2}.$
From
$\langle v,w\rangle_{\mathbb{H}}$ $=\langle v_{1}+v_{2}j,w_{1}+w_{2}j\rangle$
$=\langle v_{1},w_{1}\rangle-j\langle v_{2},w_{2}\rangle j-j\langle v_{2},w_{1}\rangle+\langle v_{1},w_{2}\rangle j,$
$=\langle v_{1},w_{1}\rangle+\langle\overline{v_{2}},\overline{w_{2}}\rangle-(\langle\overline{v_{2}},\overline{w_{1}}\rangle+\langle v_{1},-w_{2}\rangle)j,$
$=\langle[v]_{\mathbb{C}},[w]_{\mathbb{C}}\rangle_{\mathbb{C}}-\langle[v]_{\mathbb{C}},[wj]_{\mathbb{C}}\rangle_{\mathbb{C}}j$
we get the analogues of (3.22) and (3.23)
$\mathop{\rm Co}\nolimits_{1}(\langle v,w\rangle_{\mathbb{H}})=\langle[v]_{\mathbb{C}},[w]_{\mathbb{C}}\rangle_{\mathbb{C}},\qquad\mathop{\rm Co}\nolimits_{2}(\langle v,w\rangle_{\mathbb{H}})=-\langle[v]_{\mathbb{C}},[wj]_{\mathbb{C}}\rangle_{\mathbb{C}}.$ (3.31)
$\langle[v]_{\mathbb{C}},[vj]_{\mathbb{C}}\rangle_{\mathbb{C}}=0.$ (3.32)
The analogue of (3.21) for $v=z+wj$ is
$v(\alpha+\beta j)]_{\mathbb{C}}=[v\alpha+v\beta j]_{\mathbb{C}}=[v\alpha+vj\overline{\beta}]_{\mathbb{C}}=[v]_{\mathbb{C}}\alpha+[vj]_{\mathbb{C}}\overline{\beta},\qquad\alpha,\beta\in\mathbb{C},$ (3.33)
where
$v]_{\mathbb{C}}=\begin{pmatrix}z\cr\overline{w}\end{pmatrix},\quad[vj]_{\mathbb{C}}=\begin{pmatrix}-w\cr\overline{z}\end{pmatrix},\qquad\langle[v]_{\mathbb{C}},[vj]_{\mathbb{C}}\rangle_{\mathbb{C}}=0.$
Thus $\cdot$$_{\mathbb{C}}$ maps $k$-dimensional $\mathbb{H}$-subspaces of $\mathbb{H}^{d}$ to $(2k)$-dimensional $\mathbb{C}$-subspaces of $\mathbb{C}^{2d}$.
Let $L:\mathbb{H}^{n}\to\mathbb{H}^{m}$ be an $\mathbb{H}$-linear map be represented as a $\mathbb{C}$-linear map $L]_{\mathbb{C}}:\mathbb{C}^{2n}\to\mathbb{C}^{2m}$ under this identification, i.e., $L]_{\mathbb{C}}:=[\cdot]_{\mathbb{C}}L[\cdot]_{\mathbb{C}}^{-1}$. In view of (3.29), its standard matrix $L]_{\mathbb{H}}\in\mathbb{H}^{m\times n}$ has a unique decomposition
$L]_{\mathbb{H}}=A+Bj,\qquad A,B\in\mathbb{C}^{m\times n}.$
We have
$L(z+wj)$ $=(A+Bj)(z+wj)=Az+Awj+Bjz+Bjwj$
$=Az+Awj+B\overline{z}j-B\overline{w}=Az-B\overline{w}+\overline{(\overline{B}z+\overline{A}\overline{w})}j,$
and
$L^{*}]_{\mathbb{H}}=(A+Bj)^{*}=A^{*}+(-j)B^{*}=A^{*}-\overline{B^{*}}j=A^{*}-B^{T}j,$
so that
$[L]_{\mathbb{C}}$ $=\begin{pmatrix}A&-{B}\cr\overline{B}&\overline{A}\end{pmatrix},\qquad\mathop{\rm rank}\nolimits([L]_{\mathbb{C}})=2\mathop{\rm rank}\nolimits([L]_{\mathbb{H}}),$
$[L^{*}]_{\mathbb{C}}$ $=\begin{pmatrix}A^{*}&B^{T}\cr-B^{*}&A^{T}\end{pmatrix}=[L]_{\mathbb{C}}^{*}.$
The other observations for the previous case also hold (see Theorem 3.3), in particular
  • Equi-isoclinic subspaces of dimension $r$ in $\mathbb{H}^{d}$ correspond to equi-isoclinic subspaces of dimension $2r$ in $\mathbb{C}^{2d}$, and similarly for equichordal subspaces.
We now seek the analogue of Theorem 3.1, this time starting with the development in terms of the Gramian. The variational characterisation for $V=[v_{1},\ldots,v_{n}]$ being a tight frame for $\mathbb{H}^{d}$ and for $V_{\mathbb{C}}:=\bigl{[}[v_{1}]_{\mathbb{C}},\ldots,[v_{n}]_{\mathbb{C}}\bigr{]}$ being a tight frame for $\mathbb{C}^{2d}$ are
$\|V^{*}V\|_{F}^{2}={1\over d}(\mathop{\rm trace}\nolimits(V^{*}V))^{2},\qquad\|V_{\mathbb{C}}^{*}V_{\mathbb{C}}\|_{F}^{2}={1\over 2d}(\mathop{\rm trace}\nolimits(V_{\mathbb{C}}^{*}V_{\mathbb{C}}))^{2}.$
Since $\mathop{\rm trace}\nolimits(V^{*}V)=\mathop{\rm trace}\nolimits(V_{\mathbb{C}}^{*}V_{\mathbb{C}})$, a tight frame for $\mathbb{H}^{d}$ gives a tight frame for $\mathbb{C}^{2d}$ if and only if
$2\|V_{\mathbb{C}}^{*}V_{\mathbb{C}}\|_{F}^{2}-\|V^{*}V\|_{F}^{2}=0.$ (3.34)
Writing this explicitly in terms of the Gramian $V^{*}V$ gives the following.
Lemma 3.1
Let $V=[v_{1},\ldots,v_{n}]=V_{1}+V_{2}j\in\mathbb{H}^{d\times n}$. Then the following are equivalent
  1. (i)
    $V_{\mathbb{C}}=\bigl{[}[v_{1}]_{\mathbb{C}},\ldots,[v_{n}]_{\mathbb{C}}\bigr{]}=\begin{pmatrix}V_{1}\cr\overline{V_{2}}\end{pmatrix}\in\mathbb{C}^{2d\times n}$ is a tight frame for $\mathbb{C}^{2d}$.
  2. (ii)
    $V_{1}V_{1}^{*}=V_{2}V_{2}^{*}={1\over 2}AI,\qquad V_{1}V_{2}^{T}=V_{2}V_{1}^{T}=0,\qquad A:={1\over d}\sum_{j}\|v_{j}\|^{2}.$
  3. (iii)
    $(V_{1}^{*}V_{1}+V_{2}^{T}\overline{V_{2}})^{2}={1\over 2}A\,(V_{1}^{*}V_{1}+V_{2}^{T}\overline{V_{2}}),\qquad A:={1\over d}\sum_{j}\|v_{j}\|^{2}.$
  4. (iv)
    $\|\mathop{\rm Co}\nolimits_{1}(V^{*}V)\|_{F}^{2}=\sum_{j}\sum_{k}|\mathop{\rm Co}\nolimits_{1}(\langle v_{j},v_{k}\rangle)|^{2}={1\over 2d}\Bigl{(}\sum_{j}\|v_{j}\|^{2}\Bigr{)}^{2}.$
Proof: In terms of the frame operator, the condition (i) is
$\begin{pmatrix}V_{1}\cr\overline{V_{2}}\end{pmatrix}\begin{pmatrix}V_{1}\cr\overline{V_{2}}\end{pmatrix}^{*}=\begin{pmatrix}V_{1}\cr\overline{V_{2}}\end{pmatrix}\begin{pmatrix}V_{1}^{*}&V_{2}^{T}\end{pmatrix}^{*}=\begin{pmatrix}V_{1}V_{1}^{*}&V_{1}V_{2}^{T}\cr\overline{V_{2}}V_{1}^{*}&\overline{V_{2}}V_{2}^{T}\end{pmatrix}={1\over 2}A\begin{pmatrix}I&0\cr 0&I\end{pmatrix}$,
where $dA=\sum_{j}\|v_{j}\|^{2}$, which is clearly equivalent to (ii).
In terms of the Gramian $V_{\mathbb{C}}^{*}V_{\mathbb{C}}=V_{1}^{*}V_{1}+V_{2}^{T}\overline{V_{2}}$ being (a multiple of) an orthogonal projection matrix (Proposition 2.1), the condition (i) is (iii).
In terms of the variational characterisation (Theorem 2.1), the condition (i) is
$\sum_{j}\sum_{k}|\langle[v_{j}]_{\mathbb{C}},[v_{k}]_{\mathbb{C}}\rangle|^{2}={1\over 2d}\sum_{j}\Bigl{(}\|[v_{j}]_{\mathbb{C}}\|^{2}\Bigr{)}^{2},$
which can be written as (iv), since $\langle[v]_{\mathbb{C}},[w]_{\mathbb{C}}\rangle_{\mathbb{C}}=\mathop{\rm Co}\nolimits_{1}(\langle v,w\rangle_{\mathbb{H}})$ and $\|[v]_{\mathbb{C}}\|=\|v\|_{\mathbb{H}}$.
We observe that condition the (iv) depends only on $V$ up to unitary equivalence, and so the others do also.
Theorem 3.2
Let $\cdot]_{\mathbb{C}}:\mathbb{H}^{d}\to\mathbb{C}^{2d}$ be the correspondence (3.30) between $\mathbb{H}^{d}$ and $\mathbb{C}^{2d}$. Then
  1. 1.
    Tight frames for $\mathbb{C}^{2d}$ correspond to tight frames for $\mathbb{H}^{d}$.
  2. 2.
    A tight frame $V=[v_{1},\ldots,v_{n}]$ for $\mathbb{H}^{d}$ corresponds to a tight frame for $\mathbb{C}^{2d}$ if and only if it satisfies
    $\sum_{j}\sum_{k}|\mathop{\rm Co}\nolimits_{1}(\langle v_{j},v_{k}\rangle)|^{2}=\sum_{j}\sum_{k}|\mathop{\rm Co}\nolimits_{2}(\langle v_{j},v_{k}\rangle)|^{2}.$ (3.35)
Proof: The sequence $V=V_{1}+V_{2}j$ is a tight frame for $\mathbb{H}^{d}$ if and only if
$VV^{*}=(V_{1}+V_{2}j)(V_{1}^{*}-V_{2}^{T}j)=(V_{1}V_{1}^{*}+V_{2}V_{2}^{*})+(V_{2}V_{1}^{T}-V_{1}V_{2}^{T})j=AI,$
which is clearly satisfied if $V$ corresponds to a tight frame for $\mathbb{C}^{2d}$ (by Lemma 3.1).
The variational characterisation for being a tight frame for $\mathbb{H}^{d}$ and for $\mathbb{C}^{2d}$ are
$\|V^{*}V\|_{F}^{2}={1\over d}\Bigl{(}\sum_{j}\|v_{j}\|^{2}\Bigr{)}^{2},\qquad\|\mathop{\rm Co}\nolimits_{1}(V^{*}V)\|_{F}^{2}={1\over 2d}\Bigl{(}\sum_{j}\|v_{j}\|^{2}\Bigr{)}^{2}.$
Hence, if $V$ gives a tight frame for $\mathbb{H}^{d}$, then it gives a tight frame for $\mathbb{C}^{2d}$ if and only if
$2\|\mathop{\rm Co}\nolimits_{1}(V^{*}V)\|_{F}^{2}-\|V^{*}V\|_{F}^{2}=0.$
Since $|\langle v_{j},v_{k}\rangle|^{2}=|\mathop{\rm Co}\nolimits_{1}(\langle v_{j},v_{k}\rangle)|^{2}+|\mathop{\rm Co}\nolimits_{2}(\langle v_{j},v_{k}\rangle)|^{2}$, this is (3.35).
The conditions (3.28) and (3.35) can be written insightfully as
$\|\mathop{\rm Re}\nolimits(V^{*}V)\|_{F}=\|\mathop{\rm Im}\nolimits(V^{*}V)\|_{F},\qquad\|\mathop{\rm Co}\nolimits_{1}(V^{*}V)\|_{F}=\|\mathop{\rm Co}\nolimits_{2}(V^{*}V)\|_{F}.$
Example 3.2
Let $V=[1,i,j,k]$, which is a tight frame for $\mathbb{H}$. The Gramian is
$V^{*}V=\begin{pmatrix}1&i&j&k\cr-i&1&-k&j\cr-j&k&1&-i\cr-k&-j&i&1\end{pmatrix}=\begin{pmatrix}1&i&0&0\cr-i&1&0&0\cr 0&0&1&-i\cr 0&0&i&1\end{pmatrix}+\begin{pmatrix}0&0&1&i\cr 0&0&-i&1\cr-1&i&0&0\cr-i&-1&0&0\end{pmatrix}j,$
so this gives a tight frame for $\mathbb{C}^{2}$, i.e., $W=[e_{1},ie_{1},e_{2},ie_{2}]$, with Gramian
$W^{*}W=\begin{pmatrix}1&i&0&0\cr-i&1&0&0\cr 0&0&1&i\cr 0&0&-i&1\end{pmatrix}=\begin{pmatrix}1&0&0&0\cr 0&1&0&0\cr 0&0&1&0\cr 0&0&0&1\end{pmatrix}+i\begin{pmatrix}0&1&0&0\cr-1&0&0&0\cr 0&0&0&1\cr 0&0&-1&0\end{pmatrix}$,
so that this in turn gives a tight frame for $\mathbb{R}^{4}$, i.e., $e_{1},e_{3},e_{2},e_{4}]$.
Example 3.3
Consider the Gramian of the SIC of four vectors in $\mathbb{C}^{2}$ (Example 2.1). The contribution to $\|V^{*}V\|_{F}$ of the diagonal entries, which are all real, is $4$, and for the off diagonal entries it is $12{1\over 3}=4$. Thus the SIC corresponds to a tight frame for $\mathbb{R}^{4}$ if and only if its vectors can be scaled so that the off diagonal entries of the Gramian are pure imaginary. This can in fact be done, e.g., take $V=[v,iSv,i\Omega v,-S\Omega v]$, to obtain
$\begin{pmatrix}\mathop{\rm Re}\nolimits(V)\cr\mathop{\rm Im}\nolimits(V)\end{pmatrix}=\begin{pmatrix}a&-b&0&b\cr b&0&b&-a\cr 0&b&a&b\cr b&a&-b&0\end{pmatrix},\qquad a={\sqrt{3+\sqrt{3}}\over\sqrt{6}},\ b={\sqrt{3-\sqrt{3}}\over 2\sqrt{3}}.$
This is an orthonormal basis, by Proposition 2.1, or directly by using (3.22). Hence there is a norm-preserving (invertible) $\mathbb{R}$-linear map $\mathbb{C}^{2}\to\mathbb{R}^{4}$ which maps the SIC to an orthonormal basis.
We now summarise some basic results about $\cdot]_{\mathbb{F}}$, $\mathbb{F}=\mathbb{R},\mathbb{C}$, and the associated linear maps, in a unified form. We first observe that in the literature, there is some variation in the definitions, in particular, the ordering of $v]_{\mathbb{R}}$ can be either of
$v]_{\mathbb{R}}=\begin{pmatrix}\mathop{\rm Re}\nolimits(v)\cr\mathop{\rm Im}\nolimits(v)\end{pmatrix},\qquad\begin{pmatrix}\mathop{\rm Re}\nolimits(v_{1})\cr\mathop{\rm Im}\nolimits(v_{1})\cr\vdots\cr\mathop{\rm Re}\nolimits(v_{d})\cr\mathop{\rm Im}\nolimits(v_{d})\end{pmatrix}$,
and similarly for $v]_{\mathbb{C}}$. In the latter case (cf [Hog77], [Rod14] for $v]_{\mathbb{C}}$), the matrix representation $A]_{\mathbb{R}}$ is then obtained by replacing the entry $a_{jk}$ of the matrix $A$ by the matrix
$\begin{pmatrix}\mathop{\rm Re}\nolimits(a_{jk})&-\mathop{\rm Im}\nolimits(a_{jk})\cr\mathop{\rm Im}\nolimits(a_{jk})&\mathop{\rm Re}\nolimits(a_{jk})\end{pmatrix}$.
Our choice of the former was governed by the simpler formulas (cf [Coh80]). Indeed, with $L=A+iB,A+Bj$ (respectively), we have the explicit formulas
$L]_{\mathbb{F}}=\begin{pmatrix}A&-B\cr\overline{B}&\overline{A}\end{pmatrix},\qquad[L^{*}]_{\mathbb{F}}=\begin{pmatrix}A^{*}&B^{T}\cr-B^{*}&A^{T}\end{pmatrix}=[L]_{\mathbb{F}}^{*},\qquad\mathbb{F}=\mathbb{R},\mathbb{C}.$ (3.36)
Theorem 3.3
The $\mathbb{F}$-linear maps $\cdot]_{\mathbb{F}}$, $\mathbb{F}=\mathbb{R},\mathbb{C}$ given by (3.20) and (3.30) have the following properties
  1. (a)
    They map $r$-dimensional subspaces to $(2r)$-dimensional subspaces.
  2. (b)
    They preserve the Euclidean norm of a vector.
  3. (c)
    They map orthogonal vectors to orthogonal vectors.
  4. (d)
    They map tight frames satisfying (3.28) and (3.35), respectively, to tight frames.
  5. (e)
    They map equi-isoclinic $r$-subspaces to equi-isoclinic $(2r)$-subspaces.
  6. (f)
    They map equichordal $r$-subspaces to equichordal $(2r)$-subspaces.
Moreover, the associated $\mathbb{F}$-linear maps $L\mapsto[L]_{\mathbb{F}}$ to matrices over $\mathbb{F}$ satisfy
  1. (i)
    $AB]_{\mathbb{F}}=[A]_{\mathbb{F}}[B]_{\mathbb{F}}$, $\lambda A]_{\mathbb{F}}=\lambda[A]_{\mathbb{F}}$, $\lambda\in\mathbb{R}$, and $A^{*}]_{\mathbb{F}}=[A]_{\mathbb{F}}^{*}$.
  2. (ii)
    They map rank $r$ linear maps to rank $2r$ linear maps.
  3. (iii)
    They map invertible linear maps to invertible linear maps, with $A^{-1}]_{\mathbb{F}}=[A]_{\mathbb{F}}^{-1}$.
  4. (iv)
    They map self adjoint operators to self adjoint operators.
  5. (v)
    They map unitary operators to unitary operators.
  6. (vi)
    They map orthogonal projections to orthogonal projections, and in particular the identity to the identity.
Proof: For the first part, (a) has already been observed, (b) and (c) follow directly from (3.22) and (3.31), (d) follows from Theorems 3.1 and 3.2, and (e) and (f) follow from the definitions (2.19) and (2.18), and the facts (i), (ii), (vi).
Now the second part. The first part of (i) follows from the definition, and the second part was a calculation that we did in each case. For (ii), we have $\mathop{\rm ker}\nolimits([L]_{\mathbb{F}})=[\mathop{\rm ker}\nolimits(L)]_{\mathbb{F}}$, and the result follows from (a), with (iii) being a special case. If $A$ is invertible, then (i) gives $I=[I]_{\mathbb{F}}=[AA^{-1}]_{\mathbb{F}}=[A]_{\mathbb{F}}[A^{-1}]_{\mathbb{F}}$, which gives the formula for the inverse. The properties (iv), (v) and (vi) are straightforward calculations using (3.36).
Example 3.4
From the observation
$j(A_{1}+A_{2}j)=(\overline{A_{1}}+\overline{A_{2}}j)j,\qquad A_{1},A_{2}\in\mathbb{C}^{m\times n},$
it follows that the image of the $m\times n$ matrices over $\mathbb{H}$ is
$\mathbb{H}^{m\times n}]_{\mathbb{C}}=\{A\in\mathbb{C}^{2m\times 2n}:J_{m}A=\overline{A}J_{n}\},\qquad J_{\ell}:=[jI_{\ell}]_{\mathbb{C}}=\begin{pmatrix}0&-I_{\ell}\cr I_{\ell}&0\end{pmatrix}$.
Example 3.5
If $G$ is a group of $d\times d$ matrices over $\mathbb{C}$ or $\mathbb{H}$, then it follows from Theorem 3.3 that $G]_{\mathbb{F}}=\{[g]_{\mathbb{F}}:g\in G\}$ is an isomorphic group of $(2d)\times(2d)$ matrices. As an example, the quaternions $Q_{8}=\{\pm 1,\pm i,\pm j,\pm k\}$ are generated by $i$ and $j$, and so the groups of unitary matrices $Q_{8}]_{\mathbb{C}}$ and $Q_{8}]_{\mathbb{C}}]_{\mathbb{R}}$ are generated by
$i]_{\mathbb{C}}=\begin{pmatrix}i&0\cr 0&-i\end{pmatrix},\qquad[j]_{\mathbb{C}}=\begin{pmatrix}0&-1\cr 1&0\end{pmatrix}$,
$i]_{\mathbb{C}}]_{\mathbb{R}}=\begin{pmatrix}0&0&-1&0\cr 0&0&0&1\cr 1&0&0&0\cr 0&-1&0&0\end{pmatrix},\qquad[[j]_{\mathbb{C}}]_{\mathbb{R}}=\begin{pmatrix}0&-1&0&0\cr 1&0&0&0\cr 0&0&0&-1\cr 0&0&1&0\end{pmatrix}$$\qquad$
respectively. These representations of $Q_{8}$ are well known.
Example 3.6
If $V=[v_{1},\ldots,v_{n}]\in\mathbb{H}^{d\times n}$ gives a tight frame of $n$ vectors for $\mathbb{H}^{d}$, i.e., $VV^{*}=AI$, then
$V]_{\mathbb{C}}=\bigl{[}[v_{1}]_{\mathbb{C}},\ldots,[v_{n}]_{\mathbb{C}},[v_{1}j]_{\mathbb{C}},\ldots,[v_{n}j]_{\mathbb{C}}\bigr{]}$
gives a tight frame of $2n$ vectors for $\mathbb{C}^{2d}$.
The equiangular lines in $\mathbb{H}^{2}$ of [ET20] were obtained by considering equi-isoclinic planes in $\mathbb{C}^{4}$. We now explain the mechanism.
Example 3.7
Associated with a unit vector $v_{a}\in\mathbb{H}^{d}$, we have
$V_{a}:=[[v_{a}]_{\mathbb{C}},[v_{a}j]_{\mathbb{C}}]\in\mathbb{C}^{2d\times 2},$
with orthonormal columns which span a plane in $\mathbb{C}^{2d}$. The entries of the “block Gramian” for $V=[V_{1},\ldots,V_{n}]$ are $V_{a}^{*}V_{b}$ (with $V_{a}^{*}V_{a}=I$). These satisfy
$(V_{a}^{*}V_{b})^{*}(V_{a}^{*}V_{b})=\begin{pmatrix}|\langle v_{a},v_{b}\rangle_{\mathbb{H}}|^{2}&0\cr 0&|\langle v_{a},v_{b}\rangle_{\mathbb{H}}|^{2}\end{pmatrix}$, (3.37)
so that
$|\langle v_{a},v_{b}\rangle|^{2}=\lambda\quad\Longleftrightarrow\quad(V_{a}^{*}V_{b})^{*}(V_{a}^{*}V_{b})=\lambda I.$
Thus $(v_{a})$ gives a set of equiangular lines in $\mathbb{H}^{d}$ if and only if the off diagonal entries of the block Gramian $V_{1},\ldots,V_{n}]^{*}[V_{1},\ldots,V_{n}]$ are unitary matrices, up to a fixed scalar. An $n\times n$ block matrix with this structural form ($2\times 2$ blocks, positive semi-definite of rank $2d$), which corresponds to equi-isoclinic planes in $\mathbb{C}^{2d}$, can then be mapped back (under $\cdot]_{\mathbb{C}}^{-1}$) to the Gramian of $n$ equiangular lines in $\mathbb{H}^{d}$, see Theorem 13, [ET20].
The equation (3.37) follows by a direct calculation, e.g., using (3.31), we have
$(V_{b}^{*}V_{a}V_{a}^{*}V_{b})_{11}$ $=[v_{b}]_{\mathbb{C}}^{*}[v_{a}]_{\mathbb{C}}[v_{a}]_{\mathbb{C}}^{*}[v_{b}]_{\mathbb{C}}+[v_{b}]_{\mathbb{C}}^{*}[v_{a}j]_{\mathbb{C}}[v_{a}j]_{\mathbb{C}}^{*}[v_{b}]_{\mathbb{C}}$
$=\langle[v_{b}]_{\mathbb{C}},[v_{a}]_{\mathbb{C}}\rangle_{\mathbb{C}}\langle[v_{a}]_{\mathbb{C}},[v_{b}]_{\mathbb{C}}\rangle_{\mathbb{C}}+\langle[v_{b}]_{\mathbb{C}},[v_{a}j]_{\mathbb{C}}\rangle_{\mathbb{C}}\langle[v_{a}j]_{\mathbb{C}},[v_{b}]_{\mathbb{C}}\rangle_{\mathbb{C}}$
$=|\mathop{\rm Co}\nolimits_{1}(\langle v_{a},v_{b}\rangle_{\mathbb{H}})|^{2}+|\mathop{\rm Co}\nolimits_{2}(\langle v_{a},v_{b}\rangle_{\mathbb{H}})|^{2}=|\langle v_{a},v_{b}\rangle_{\mathbb{H}}|^{2},$
$(V_{b}^{*}V_{a}V_{a}^{*}V_{b})_{12}$ $=[v_{b}]_{\mathbb{C}}^{*}[v_{a}]_{\mathbb{C}}[v_{a}]_{\mathbb{C}}^{*}[v_{b}j]_{\mathbb{C}}+[v_{b}]_{\mathbb{C}}^{*}[v_{a}j]_{\mathbb{C}}[v_{a}j]_{\mathbb{C}}^{*}[v_{b}j]_{\mathbb{C}}$
$=\langle[v_{b}]_{\mathbb{C}},[v_{a}]_{\mathbb{C}}\rangle_{\mathbb{C}}\langle[v_{a}]_{\mathbb{C}},[v_{b}j]_{\mathbb{C}}\rangle_{\mathbb{C}}+\langle[v_{b}]_{\mathbb{C}},[v_{a}j]_{\mathbb{C}}\rangle_{\mathbb{C}}\langle[v_{a}j]_{\mathbb{C}},[v_{b}j]_{\mathbb{C}}\rangle_{\mathbb{C}}$
$=\mathop{\rm Co}\nolimits_{1}(\langle v_{b},v_{a}\rangle_{\mathbb{H}})(-\mathop{\rm Co}\nolimits_{2}(\langle v_{a},v_{b}\rangle_{\mathbb{H}}))-\mathop{\rm Co}\nolimits_{2}(\langle v_{b},v_{a}\rangle_{\mathbb{H}})\mathop{\rm Co}\nolimits_{1}(\langle v_{a}j,v_{b}j\rangle_{\mathbb{H}})$
$=-\mathop{\rm Co}\nolimits_{1}(\langle v_{b},v_{a}\rangle_{\mathbb{H}})\mathop{\rm Co}\nolimits_{2}(\langle v_{a},v_{b}\rangle_{\mathbb{H}})+\mathop{\rm Co}\nolimits_{2}(\langle v_{b},v_{a}\rangle_{\mathbb{H}})\mathop{\rm Co}\nolimits_{1}(\langle v_{b},v_{a}\rangle_{\mathbb{H}})=0,$
where in the second to last equality we used $\mathop{\rm Co}\nolimits_{2}(\overline{q})=-\mathop{\rm Co}\nolimits_{2}(q)$, $q\in\mathbb{H}$.
Here is a construction of equiangular lines going in the opposite direction.
Example 3.8
We consider the construction of $64$ equiangular lines in $\mathbb{C}^{8}$ by [Hog98]. These were obtained by finding $64$ unit vectors in $\mathbb{H}^{4}$ with angles ${1\over 9},{1\over 3}$ (as vertices of a quaternionic polytope). These were then mapped by $\cdot$$_{\mathbb{C}}$ to $64$ equiangular vectors in $\mathbb{C}^{8}$. We note that for $v,w\in\mathbb{H}^{d}$, $\alpha\in\mathbb{H}$, (3.31) gives
$\langle[v\alpha]_{\mathbb{C}},[w]_{\mathbb{C}}\rangle_{\mathbb{C}}$ $=\mathop{\rm Co}\nolimits_{1}(\langle v\alpha,w\rangle_{\mathbb{H}})=\mathop{\rm Co}\nolimits_{1}(\langle v,w\rangle_{\mathbb{H}}\alpha)$
$=\mathop{\rm Co}\nolimits_{1}(\alpha)\mathop{\rm Co}\nolimits_{1}(\langle v,w\rangle_{\mathbb{H}})-\mathop{\rm Co}\nolimits_{2}(\alpha)\overline{\mathop{\rm Co}\nolimits_{2}(\langle v,w\rangle_{\mathbb{H}})},$
so that multiplying vectors in $\mathbb{H}^{d}$ by noncomplex unit scalars in $\mathbb{H}$ can change the angle between their images in $\mathbb{C}^{2d}$.
4 Group frames and $G$-matrices
Many tight frames of interest are the orbit of one or more vectors under the unitary action of a group, e.g., the Weyl-Heisenberg SICs. There is a well developed theory of such group frames based in the theory of group representations (over $\mathbb{R}$ and $\mathbb{C}$) [VW05], [Wal13], [VW16], [Wal18]. We now give an indication of how this theory extends to representations over $\mathbb{H}$ (see [SS95]).
A representation of a finite abstract group $G$ on $\mathbb{H}^{d}$ is a group homomorphism $\rho:G\to\mathop{\it GL}\nolimits(\mathbb{H}^{d})$ from $G$ to the invertible $d\times d$ matrices over $\mathbb{H}$, with equivalence defined in the usual way. We will consider only unitary representations, i.e.. those where the matrices $\rho(g)$ are unitary. For these, we will write the unitary action as $gv:=\rho(g)v$, and we note that $g^{*}v=g^{-1}v$. A frame (sequence of vectors) of the form $(gv)_{g\in G}$ is said to be a group frame (or $G$-frame) [Wal20a]. The frame operator of a group frame $(gv)_{g\in G}$ commutes with the frame operator, i.e.,
$S(hv)=\sum_{g\in G}gv\langle gv,hv\rangle=h\sum_{g\in G}h^{-1}gv\langle h^{-1}gv,v\rangle=hS(v),\quad h\in G,\ v\in\mathbb{H}^{d}.$ (4.38)
The Gramian of a group matrix has entries of the form
$\langle hv,gv\rangle=\langle g^{-1}hv,v\rangle.$
A matrix $A=[a_{gh}]_{g,h\in G}\in\mathbb{H}^{G\times G}$ is a $G$-matrix (or group matrix) if there exists a function $\nu:G\to\mathbb{H}$ such that
$a_{gh}=\nu(g^{-1}h),\qquad\forall g,h\in G.$
The Gramian of a $G$-frame is a $G$-matrix, and conversely if the Gramian of a frame $(v_{g})_{g\in G}$ with vectors indexed by $G$ is a $G$-matrix, then it is a $G$-frame (adapt the proof of [Wal18] Theorem 10.3). An action (representation) of $G$ on $\mathbb{H}^{d}$ is irreducible if the only $G$-invariant subspaces of $\mathbb{H}^{d}$ are 0 and $\mathbb{H}^{d}$, i.e., $\mathop{\rm span}\nolimits_{\mathbb{H}}\{gv\}_{g\in G}=\mathbb{H}^{d}$, for all $v\neq 0$.
The theory of $G$-frames for real and complex actions begins with irreducible actions, where it takes its simplest form. This extends without issue.
Proposition 4.1
Suppose that a unitary action of a group $G$ on $\mathbb{H}^{d}$ is irreducible. Then $(gv)_{g\in G}$ is a tight $G$-frame for $\mathbb{H}^{d}$ for any $v\neq 0$, i.e.,
$x={d\over|G|}{1\over\|v\|^{2}}\sum_{g\in G}gv\langle gv,x\rangle,\qquad\forall x\in\mathbb{H}^{d}.$
Proof: Fix $v\neq 0$, and let $S$ be the frame operator of $(gv)_{g\in G}$. Since $S$ is nonzero and positive semidefinite, it has an eigenvalue $\lambda>0$, with corresponding eigenvector $w$. By (4.38), $S$ commutes with the action of $g\in G$, so that
$S(gw)=g(Sw)=g(w\lambda)=(gw)\lambda,$
so that $gw$ is an eigenvector for $\lambda$. But $(gw)_{g\in G}$ spans $\mathbb{H}^{d}$, so that $S=\lambda I$, i.e., $(gv)_{g\in G}$ is a tight frame. Since $S$ is Hermitian, taking the trace gives
$\mathop{\rm trace}\nolimits(S)=\mathop{\rm Re}\nolimits(\mathop{\rm trace}\nolimits(S))=\sum_{g}\|gv\|^{2}=|G|\,\|v\|^{2}=\mathop{\rm trace}\nolimits(\lambda I)=d\lambda,$
which gives the value of $\lambda$.
The general theory [VW16], [Wal18], which allows for multiple orbits, involves the decomposition of the vector space into irreducible $G$-invariant subspaces.
Example 4.1
Each finite subgroup of $\mathbb{H}^{*}$ corresponds to a (faithful) irreducible action on $\mathbb{H}^{1}$. These subgroups were classified by Stringham [Str81]. They are the infinite families of cyclic groups (generated by the $n$-th roots of unity) and binary dihedral groups, together with the binary tetrahedral, octahedral and icosahedral groups.
Example 4.2
The group generated by the matrices
$\begin{pmatrix}0&1\cr 1&0\end{pmatrix},\quad\begin{pmatrix}1&0\cr 0&i\end{pmatrix},\quad\begin{pmatrix}1&0\cr 0&j\end{pmatrix}$,
has an irreducible unitary action on $\mathbb{H}^{2}$. It consists of all $128$ invertible matrices with two zero entries and two entries in $Q_{8}$. It contains the scalar matrices from $Q_{8}$ and its center is $\pm I$. Thus each orbit can be viewed as $16$ lines in $\mathbb{H}^{2}$ (as a left vector space). This is an example of a (quaternionic) reflection group, i.e., a finite group generated by reflections (linear maps which act as the identity on a hyperplane). The finite irreducible quaternionic reflection groups have been classified (up to conjugacy) by Cohen [Coh80].
It is expected that the highly symmetric tight frames of [BW13] corresponding to complex reflection groups could be extended to the quaternionic reflection groups. In this regard we note the regular quaternionic polytopes have been classified by [Cuy95].
For $G$ abelian, there are a finite number of tight $G$-frames (called harmonic frames) that can be obtained by “taking rows of the character table” (see [VW05], [CW11]). We now give an example to show how this can be extended to the quaternionic setting.
Example 4.3
(Quaternionic harmonic frames). The irreducible representations over $\mathbb{C}$ for an abelian group $G$ are all one-dimensional (this characterises abelian groups), and these “rows” of the character table are orthogonal, so by taking a set of rows of the character table one obtains a tight $G$-frame. Consider the quaternion group $G=Q_{8}$. This has four $1$-dimensional and one $2$-dimensional irreducible representations over $\mathbb{C}$. The $2$-dimensional absolutely irreducible representation splits into four $1$-dimensional representations over $\mathbb{H}$, corresponding to the outer automorphisms of the quaternions. In this way, one obtains a character table
$\begin{array}[]{lcccccccc}q\in Q_{8}&1&-1&i&-i&j&-j&k&-k\\ \hline\cr\chi_{1}&1&1&1&1&1&1&1&1\\ \chi_{2}&1&1&1&1&-1&-1&-1&-1\\ \chi_{3}&1&1&-1&-1&1&1&-1&-1\\ \chi_{4}&1&1&-1&-1&-1&-1&1&1\\ \chi_{5}&1&-1&i&-i&j&-j&k&-k\\ \chi_{6}&1&-1&j&-j&i&-i&-k&k\\ \chi_{7}&1&-1&-i&i&k&-k&j&-j\\ \chi_{8}&1&-1&k&-k&-i&i&-j&j\end{array}$
where the rows are orthogonal (cf [SS95]). Taking rows gives a $G$-frame. The columns of the character table are also orthogonal, so taking columns also gives a tight frame, but these are not $G$-frames, in general (as follows for abelian groups by Pontryagin duality). As an example, the frame obtained by taking the characters $\chi_{1}$ and $\chi_{5}$ (rows $1$ and $5$ of the character table) gives a (unit-norm) tight $Q_{8}$-frame for $\mathbb{H}^{2}$, with the inner products $\{1\pm 1,1\pm i,1\pm j,1\pm k\}$ occurring exactly once in every row (column) of the Gramian. This frame has two angles: each vector is orthogonal to one other, and makes a fixed angle with all the others.
5 Projective unitary equivalence
Finally, we consider the equivalence of vectors thought of as lines in $\mathbb{H}^{d}$. Here the noncommutativity of scalar multiplication considerably complicates the theory.
We say that two sequences of vectors $(v_{j})$ and $(w_{j})$ in $\mathbb{H}^{d}$ are projectively unitarily equivalent if there exists a unitary $U$ and unit norm scalars $\alpha_{j}$ with
$w_{j}=(Uv_{j})\alpha_{j},\qquad\forall j.$
Clearly, projective unitary equivalence is an equivalence relation. Moreover, one can define a projective unitary symmetry group of $(v_{j})_{j\in J}$ to be all the permutations $\sigma:J\to J$ for which $(v_{j})$ and $(v_{\sigma j})$ are projectively unitarily equivalent (cf [CW18]).
To make a workable theory, one now needs a way to recognise projective unitary equivalence. In terms of the Gramians $V=[v_{j}]$ and $W=[w_{j}]$, the formal definition says that
$W^{*}W=C^{*}V^{*}U^{*}UVC=C^{*}V^{*}VC,\qquad C:=\mathop{\rm diag}\nolimits(\alpha_{j}),$
i.e.,
$\langle w_{j},w_{k}\rangle=\overline{\alpha_{k}}\langle v_{j},v_{k}\rangle\alpha_{j}.$ (5.39)
This leads to a “linear system” $C(W^{*}W)=(V^{*}V)C$ in the scalars $\alpha_{j}$. However, due to the noncommutativity of the quaternions, this can not be solved by Gauss elimination, unless one first converts it to a linear system over $\mathbb{R}$ (in the coordinates of the $\alpha_{j}$). What is usually done in the real and complex cases is to consider a collection of invariants: the $m$-products, which completely characterise projective unitary equivalence [CW16]. We now look at the analogue of these (also see [KMW25] for subspaces of $\mathbb{C}^{d}$).
For a sequence of vectors $(v_{j})$ in $\mathbb{H}^{d}$ the $m$-products are
$\Delta(v_{j_{1}},v_{j_{2}},\ldots,v_{j_{m}}):=\langle v_{j_{1}},v_{j_{2}}\rangle\langle v_{j_{2}},v_{j_{3}}\rangle\langle v_{j_{3}},v_{j_{4}}\rangle\cdots\langle v_{j_{m}},v_{j_{1}}\rangle\in\mathbb{H}.$
The $1$-products and $2$-products are clearly projective unitary invariants, since
$\Delta(v_{j})=\|v_{j}\|^{2},\qquad\Delta(v_{j},v_{k})=|\langle v_{j},v_{k}\rangle|^{2}.$
From these, we can define the frame graph of $(v_{j})$ to be the graph with vertices $\{v_{j}\}$ and an edge between $v_{j}$ and $v_{k}$ ($j\neq k$) if and only if $\langle v_{j},v_{k}\rangle\neq 0$.
Further, since
$\Delta\bigl{(}(U$ $v_{j_{1}})\alpha_{j_{1}},(Uv_{j_{2}})\alpha_{j_{2}},\ldots,(Uv_{j_{m}})\alpha_{j_{m}}\bigr{)}$
$=\langle(Uv_{j_{1}})\alpha_{j_{1}},(Uv_{j_{2}})\alpha_{j_{2}}\rangle\langle(Uv_{j_{2}})\alpha_{j_{2}},(Uv_{j_{3}})\alpha_{j_{3}}\rangle\cdots\langle(Uv_{j_{m}})\alpha_{j_{m}},(Uv_{j_{1}})\alpha_{j_{1}}\rangle$
$=\overline{\alpha_{j_{1}}}\langle v_{j_{1}},v_{j_{2}}\rangle\alpha_{j_{2}}\overline{\alpha_{j_{2}}}\langle v_{j_{2}},v_{j_{3}}\rangle\alpha_{j_{3}}\cdots\overline{\alpha_{j_{m}}}\langle v_{j_{m}},v_{j_{1}}\rangle\alpha_{j_{1}}$
$=\overline{\alpha_{j_{1}}}\Delta(v_{j_{1}},v_{j_{2}},\ldots,v_{j_{m}})\alpha_{j_{1}}$
$=\alpha_{j_{1}}^{-1}\Delta(v_{j_{1}},v_{j_{2}},\ldots,v_{j_{m}})\alpha_{j_{1}},$
the $m$-products are projective unitary invariants of $(v_{j})$ up to similarity (congruence), and real frames are characterised by having real $m$-products. Since a quaternion $q$ is determined up to similarity (conjugation in $\mathbb{H}\setminus\{0\}$, which is equivalent to conjugation by unit scalars) by its real part $\mathop{\rm Re}\nolimits(q)$ and its norm $|q|$, we can define (reduced) $m$-products as a pair of real numbers
$\Delta_{r}(v_{j_{1}},v_{j_{2}},\ldots,v_{j_{m}}):=(\mathop{\rm Re}\nolimits(q),|q|),\qquad q=\Delta(v_{j_{1}},v_{j_{2}},\ldots,v_{j_{m}}).$
These are projective unitary invariants. For the complex case, the $m$-products are projective unitary invariants, which depend only on the cycle $(j_{1},\ldots,j_{m})$, and a small set of $m$-products corresponding to a basis for the cycle space of the frame graph of $(v_{j})$ provide a set of invariants which characterise projective unitary equivalence (see [CW16]). We can not yet make a similar claim in the quaternionic case, though we do imagine that the $m$-products do characterise projective unitary equivalence.
The dependence of $m$-products on only the associated $m$-cycle in the frame graph does follow, by the calculation (for nonzero $m$-products)
$a^{-1}\Delta(v_{j_{1}},v_{j_{2}},\ldots,v_{j_{m}})a=\Delta(v_{j_{2}},v_{j_{3}},\ldots,v_{j_{m}},v_{j_{1}}),\qquad a={\langle v_{j_{1}},v_{j_{2}}\rangle\over|\langle v_{j_{1}},v_{j_{2}}\rangle|},$
and so, in addition to the $1$-products and $2$-products, we need only consider the $m$-products for $m\geq 3$ which correspond to $m$-cycles in the frame graph, i.e., are nonzero. To check that the $m$-products for two sequences are equal (up to conjugation), it suffices to consider only the $m$-products corresponding to a cycle basis for the cycle space of the (common) frame graph:
Lemma 5.1
(Cycle decomposition) For $1\leq k\leq m$, $n\geq 1$, we have
$\Delta(v_{k},v_{k+1},\ldots,v_{m},v_{1},\ldots,v_{k-1})\Delta(v_{k},\ldots,v_{1},w_{1},\ldots,w_{n})$
$\qquad=|\langle v_{1},v_{2}\rangle|^{2}|\langle v_{2},v_{3}\rangle|^{2}\cdots|\langle v_{k-1},v_{k}\rangle|^{2}\Delta(v_{k},v_{k+1},\ldots,v_{m},v_{1},w_{1},w_{2},\ldots,w_{n}).$
Proof: Expanding the left hand side gives
$\langle v_{k},v_{k+1}\rangle\langle v_{k+1},v_{k+2}\rangle\cdots\langle v_{m-1},v_{m}\rangle\langle v_{m},v_{1}\rangle\langle v_{1},v_{2}\rangle\cdots\langle v_{k-2},v_{k-1}\rangle\langle v_{k-1},v_{k}\rangle$
$\qquad\times\langle v_{k},v_{k-1}\rangle\langle v_{k-1},v_{k-2}\rangle\cdots\langle v_{2},v_{1}\rangle\langle v_{1},w_{1}\rangle\langle w_{1},w_{2}\rangle\cdots\langle w_{n-1},w_{n}\rangle\langle w_{n},v_{k}\rangle,$
which simplifies to the right hand side, since $\langle v_{j-1},v_{j}\rangle\langle v_{j},v_{j-1}\rangle=|\langle v_{j-1},v_{j}\rangle|^{2}\in\mathbb{R}$ commutes with any quaternion.
This gives the following condition for projective unitary equivalence.
Theorem 5.1
A necessary condition for sequences $(v_{j})$ and $(w_{j})$ of $n$ vectors in $\mathbb{H}^{d}$ to be projectively unitarily equivalent is that the $m$-products corresponding to a cycle basis for the frame graph are are equal (up to conjugation).
In the complex setting, this says that the $m$-products are equal, and the converse is proved by explicitly constructing scalars $\alpha_{j}$ which satisfy (5.39). The difficulties in extending this converse to the quaternionic setting include the fact that for $w_{j}=(Uv_{j})\alpha_{j}$,
$\Delta(w_{j_{1}},w_{j_{2}},\ldots,w_{j_{m}})=\overline{\alpha_{j_{1}}}\Delta(v_{j_{1}},v_{j_{2}},\ldots,v_{j_{m}})\alpha_{j_{1}},$ (5.40)
which puts further constraints on the $\alpha_{j}$ (for $m\geq 3$ and the $m$-product nonzero). Indeed, in the complex setting one can assume that any $\alpha_{j}$ is $1$, simply by replacing $U$ by the unitary matrix $\alpha_{j}U$. Nevertheless, those parts of the theory that we do have allow us to investigate such things as the symmetries of lines, as our final example shows.
Example 5.1
Consider the six tight equiangular lines in $\mathbb{H}^{2}$ at angle $\lambda=c^{2}={2\over 5}$ of [ET20]
$v_{1}=\begin{pmatrix}1\cr 0\end{pmatrix},\ v_{2}=\begin{pmatrix}{\sqrt{2}\over\sqrt{5}}\cr{\sqrt{3}\over\sqrt{5}}\end{pmatrix},\ v_{3}=\begin{pmatrix}{\sqrt{2}\over\sqrt{5}}\cr-{\sqrt{3}\over 4\sqrt{5}}+{3\over 4}i\end{pmatrix},\ v_{4}=\begin{pmatrix}{\sqrt{2}\over\sqrt{5}}\cr-{\sqrt{3}\over 4\sqrt{5}}-{1\over 4}i+{1\over\sqrt{2}}j\end{pmatrix}$,
$v_{5}=\begin{pmatrix}{\sqrt{2}\over\sqrt{5}}\cr-{\sqrt{3}\over 4\sqrt{5}}-{1\over 4}i-{1\over 2\sqrt{2}}j+{\sqrt{3}\over 2\sqrt{2}}k\end{pmatrix},\ v_{6}=\begin{pmatrix}{\sqrt{2}\over\sqrt{5}}\cr-{\sqrt{3}\over 4\sqrt{5}}-{1\over 4}i-{1\over 2\sqrt{2}}j-{\sqrt{3}\over 2\sqrt{2}}k\end{pmatrix}$,
which are said to have “symmetry group” $A_{6}$. The reduced $m$-products $\Delta_{r}(v_{j_{1}},\ldots,v_{j_{m}})$ of distinct vectors for $m=1,2,3,4,6$ are all equal, taking the values
$(1,1),\quad({2\over 5},c^{2}),\quad({1\over 10},c^{3}),\quad(-{1\over 50},c^{4}),\quad(-{11\over 250},c^{6})$
respectively, which puts no restriction on the possible projective symmetry group of the lines. However, the reduced $5$-products (of distinct vectors) take two values
$(-{25\pm 9\sqrt{5}\over 500},c^{5}),$
and the permutations of the vectors which preserve these $5$-products is indeed $A_{6}$. Thus the projective symmetry group is a subgroup of $A_{6}$. With the present theory, this does not yet establish that $A_{6}$ is the projective symmetry group.
We now seek a corresponding projective unitary symmetry for each $\sigma\in A_{6}$, i.e., a unitary matrix $U_{\sigma}$ and corresponding scalars $\alpha_{j}$ (also depending on $\sigma$) for which
$w_{j}=v_{\sigma j}=(U_{\sigma}v_{j})\alpha_{j},\qquad\forall j.$
Once the unit scalars $\alpha_{j}$ corresponding to a basis $v_{j}]_{j\in J}$ of vectors from $(v_{j})$ are known, the matrix $U_{\sigma}$ is uniquely determined by
$U_{\sigma}[v_{j}\alpha_{j}]_{j\in J}=[v_{\sigma j}]_{j\in J}\quad\Longrightarrow\quad U_{\sigma}=[v_{\sigma j}]_{j\in J}[v_{j}\alpha_{j}]_{j\in J}^{-1},$
and it can then be checked whether or not the $U_{\sigma}$ is unitary and permutes the other lines. By (5.40), for $j,k,\ell$ distinct, the unit scalar $\alpha_{j}$ satisfies
$\alpha_{j}\Delta(v_{\sigma j},v_{\sigma k},v_{\sigma\ell})=\Delta(v_{j},v_{k},v_{\ell})\alpha_{j},$
which gives a homogeneous linear system of four equations for the four real coordinates of $\alpha_{j}$. In the cases considered, this had a unique solution of unit norm up to a choice of sign, which was made in order to obtain a unitary matrix $U_{\sigma}$. For the generators
$a=(12)(34)\quad\hbox{(order $2$)},\qquad b=(1235)(46)\quad\hbox{(order $4$)}$
for $A_{6}$, we obtained
$\alpha_{1}=\alpha_{2}=-{\sqrt{2}\over\sqrt{3}}i+{1\over\sqrt{3}}k,\qquad U_{a}=\begin{pmatrix}{2\over\sqrt{15}}i-{\sqrt{2}\over\sqrt{15}}j&{\sqrt{2}\over\sqrt{5}}i-{1\over\sqrt{5}}j\cr{\sqrt{2}\over\sqrt{5}}i-{1\over\sqrt{5}}j&-{2\over\sqrt{15}}i+{\sqrt{2}\over\sqrt{15}}j\end{pmatrix},\quad U_{a}^{2}=-I,$
and
$\alpha_{1}={1\over 2\sqrt{2}}+{\sqrt{5}\over 2\sqrt{6}}i-{3-\sqrt{5}\over 4\sqrt{3}}j-{\sqrt{5}+1\over 4}k,\quad\alpha_{2}={\sqrt{5}\over 4}+{1\over 4\sqrt{3}}i-{3\sqrt{5}+1\over 4\sqrt{6}}j-{\sqrt{5}-1\over 4\sqrt{2}},$
$U_{b}=\begin{pmatrix}{1\over 2\sqrt{5}}+{1\over 2\sqrt{3}}i+{3-\sqrt{5}\over 2\sqrt{30}}j+{\sqrt{5}+1\over 2\sqrt{10}}k&{\sqrt{3}\over 2\sqrt{10}}-{1\over 2\sqrt{2}}i+{3+\sqrt{5}\over 4\sqrt{5}}j-{\sqrt{3}\over 5+\sqrt{5}}k\cr{\sqrt{3}\over 2\sqrt{10}}+{1\over 2\sqrt{2}}i+{3-\sqrt{5}\over 4\sqrt{5}}j+{\sqrt{3}\over 5-\sqrt{5}}k&-{1\over 2\sqrt{5}}+{1\over 2\sqrt{3}}i-{3\sqrt{5}+5\over 10\sqrt{6}}j+{\sqrt{5}-1\over 2\sqrt{10}}k\end{pmatrix},\quad U_{b}^{4}=-I.$
These unitary matrices $U_{a}$ and $U_{b}$ do give the projective unitary symmetries supposed. Moreover, they generate the double cover $2\cdot A_{6}$ of $A_{6}$, and so we have verified that $A_{6}$ is indeed the projective symmetry group of the six equiangular lines in $\mathbb{H}^{2}$. We note that our method did not require prior knowledge of what the symmetry group was.
The action group of the faithful representation of $2\cdot A_{6}$ obtained in Example 5.1 contains $40$ reflections (of order $3$), and it is an irreducible reflection group which appears on the list of [Coh80]. The vectors giving the lines are eigenvectors of nontrivial elements of the group, and so the six equiangular lines in $\mathbb{H}^{2}$ can be constructed directly from the reflection group as a group frame (or even from the abstract group $2\cdot A_{6}$) [Wal24].
The sets of five and six equiangular lines in $\mathbb{H}^{2}$ were first calculated in [KF08] using the Hopf map. Though this technique does not immediately generalise to other dimensions, like that of [ET20], we recount the essential details, as it sheds further light on the geometry of these lines. The Hopf map $\psi$ maps a point $\vec{a}=(a_{1},\ldots,a_{5})$ on the unit sphere in $\mathbb{R}^{5}$ to a line in the projective space $\mathbb{H}\mathbb{P}^{1}$, i.e., a the unit vector $v\in\mathbb{H}^{2}$ in the line with $v_{2}\geq 0$, and is given by $\psi(0,0,0,0,1):=(1,0)$ and
$\psi(\vec{a}):=\begin{pmatrix}{a\over\sqrt{2(1-a_{5})}}\cr{\sqrt{1-a_{5}}\over\sqrt{2}}\end{pmatrix},\qquad a:=a_{1}+a_{2}i+a_{3}j+a_{4}k,\quad a_{5}\neq 1.$
A calculation shows that
$|\langle\psi(\vec{a}),\psi(\vec{b})\rangle_{\mathbb{H}}|^{2}$ $={1+\langle\vec{a},\vec{b}\rangle_{\mathbb{R}}\over 2},\qquad\forall\vec{a},\vec{b},$ (5.41)
so the $n\geq 3$ unit vectors $(v_{j})$, $v_{j}=\psi(\vec{v_{j}})\in\mathbb{H}^{2}$, give tight equiangular lines if and only if
$|\langle v_{j},v_{k}\rangle|^{2}={1+\langle\vec{v_{j}},\vec{v_{k}}\rangle\over 2}={n-2\over 2(n-1)}\quad\Longleftrightarrow\quad\langle\vec{v_{j}},\vec{v_{k}}\rangle=-{1\over n-1}.$
This latter condition says that the vectors $(\vec{v_{j}})$ are the vertices of a regular $n$-vertex simplex embedded in the unit sphere in $\mathbb{R}^{5}$, which can be done for $n=3,4,5,6$, with the corresponding image $(v_{j})$ giving $n$ tight equiangular lines in $\mathbb{H}^{2}$. Moreover, for $n=3$ we get real lines by choosing the simplex in $\{x:x=(x_{1},0,0,0,x_{5})\}$, and complex lines for $n=4$ by choosing the simplex in $\{x:x=(x_{1},x_{2},0,0,x_{5})\}$.
5.1 Concluding remarks
We have shown how much of the theory of tight frames extends to quaternionic Hilbert space, with the characterisation of projective unitary equivalence of frames being the aspect that most depends intrinsically on the commutativity of the complex numbers. The notions of canonical coordinates and the canonical Gramian [Wal18] also extend to $\mathbb{H}$-vector spaces. In particular, there is a unique $\mathbb{H}$-inner product for which a finite spanning set for an $\mathbb{H}$-vector space becomes a normalised tight frame.
Our focus has been on group frames and equiangular lines. The maximal set of six equiangular lines in $\mathbb{H}^{2}$ comes as the orbit of a quaternionic reflection group, just as the SIC of four equiangular lines in $\mathbb{C}^{2}$ is the orbit of a complex reflection group. However, the known SICs in $\mathbb{C}^{d}$ (with one exception) are orbits of the Weyl-Heisenberg group, which is not a reflection group for $d\geq 3$. The key to constructing quaternionic equiangular lines in this way will be knowing “the right group”. This group might come from numerical constructions, using the techniques of this last section, or from the theory of group representations over $\mathbb{H}$ (which is in its infancy). The construction of sets of tight quaternionic lines may also offer insight into Zauner’s conjecture. Another direction of similar interest is that of optimal packings in quaternionic projective space $\mathbb{H}\mathbb{P}^{k}$.
Many of our results say, in some sense, that “there is more room in $\mathbb{H}^{d}$ than in $\mathbb{C}^{d}$”. In particular, we offer the following variation of Conjecture 1, which does not implicitly reference Zauner’s conjecture.
Conjecture 2
The maximal number of quaternionic equiangular lines in $\mathbb{H}^{d}$ is strictly larger than the maximal number of complex equiangular lines in $\mathbb{C}^{d}$, for each $d\geq 2$.
Conflict of interest statement
The author declared that he has no conflict of interest.